### v4

# GLOBALIZATION? TRADE WAR? A COUNTERBALANCE PERSPECTIVE

Abstract: The embrace of globalization and protectionism among economies has ebbed and flowed over the past few decades. These fluctuations call for quantitative analytics to help countries improve their trade policies. Changing attitudes toward globalization also imply that the best trade policies vary over time and are country-specific. We argue that the imports and exports of all economies constitute a counterbalanced network where conflict and cooperation are two sides of the same coin. Quantitative competitiveness is then formulated for each country using a network counterbalance equilibrium. A country could improve its relative strength in the network by embracing globalization, protectionism, trade collaboration, or conflict. This paper presents some necessary conditions for globalization and trade wars, evaluates their side effects, derives national bargaining powers, identifies appropriate targets for conflict or collaboration, and recommends fair resolutions for trade conflicts. Data and events from 2000 to 2019 support these conditions.
Keywords:  globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma
JEL Code: C71, C78, F11, F13, F15, O24
1 Introduction
De-globalization and a trade war are two levels of economic conflict; the latter directly involves only two trade partners, whereas the former represents one country’s disintegration from the global economy. Nevertheless, both could adversely affect global economic activity. This paper applies a unified game-theoretic approach to formulate simple but decisive rules to mitigate either type of conflict. We focus on import and export activities, ignoring political, territorial, ideological, military, and cultural conflicts. However, geopolitical data have already been incorporated into trade behaviors and volumes. From this perspective, a country’s primary purpose in engaging in trade wars or globalization is to improve its competitiveness and trade balance in the trade system. Also, the engagement affects the future geopolitical landscape.
To analytically examine trade wars and globalization, we face several daunting challenges. First, there are always conflicts and cooperation in international trade. While the trade of goods and services results from bilateral collaboration, a country’s growing strength could weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study could be too impotent to capture the complexity of trade relationships. The second challenge is adequately constructing a high-level description of the interactive trade system because exports and imports have already reflected the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility or social welfare functions in this conflicting and cooperative context. Next, policy variables, such as imports and exports as production’s percentages, should be linked to these economic objectives. So, countries could apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, it is fruitful to bring data into the analytics to explain past events and find optimal solutions for the near future. The solutions could identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization. In the literature, studies on globalization often capitalize on econometric analysis, which uses historical data to estimate unknown parameters. However, the estimated effects could hardly catch up with the recent changes or seize both cooperation and non-cooperation.
We measure the relative strengths of the countries in the network by the solution to the counterbalance equilibrium, defined by Hu and Shapley, (2003), and use it as an objective in our analysis. The solution, called authority distribution, integrates both cooperation and competition inside the global economic system. The data-driven distribution is a social welfare function and an endogenously weighted average of utility functions. Moreover, it measures the competitive advantages of the countries and their long-term influences on the whole system. Also, this competitiveness brings extra welfare beyond the economy of scale and comparative advantage. The additional advantage, called the Matthew effect, is another driving force for competition after comparative advantages.
This paper attempts to provide strategic solutions for policymakers and ignores how an individual firm or economic agent improves firm’s competitiveness through research and development. Maximizing national competitive strength is a purview of government policymakers rather than trade companies, which often capitalize on comparative advantages. A country is not merely an aggregate of firms, and the interests of trade companies only sometimes align with those of policymakers, who also consider the side effects of trade, such as the shrinkage of domestic labor markets and the overseas shift of industries. So, our studies neglect the firm-based micro foundation and maximize a social welfare function, pivoting domestic production decision-making. However, the competitiveness measure is also a weighted average of comparative advantages that considerably reflect the profits of domestic firms engaging in international trade and the welfare of consumers.
When starting a trade war against a trade partner, a country raises tariffs or creates other barriers to increase competitiveness. However, the result depends on the counterpart’s reaction. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in a conflict or cooperation. On the other hand, the reaction may also benchmark their production ratio to avoid trade deficits. Using the threshold and the benchmark ratios, we derive a Nash bargaining solution for countries with heterogeneous preferences between competitiveness and trade balance, and the solution generates consistent national bargaining power. It provides a solution to resolve bilateral trade frictions.
A country with no specific targets to confront could choose between further globalization or protectionism. By minor perturbations in its overall exports, we calibrate the effect on its competitiveness from the counterbalance equilibrium. The sign of the effect tells if the country should go for more globalization or not. There are many debates over economic globalization in the literature, and Stiglitz (2002, 2017) documented their detailed viewpoints. Different viewpoints, assumptions, and considerations could reach opposite conclusions about globalization for a country. Our data-driven approach ignores arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time. By analyzing the ComTrade data of the United Nations (2021) for the years 2000 through 2019, we find that protectionism could have been a better choice for the USA and the United Kingdom but a worse option for China, Germany, Japan, and Russia if they targeted competitiveness and trade balances only. Data from 2017 also show that trade wars could be an option for the USA to maintain its competitiveness. However, adverse side effects on third parties and allies were also tangible. According to our theorem of impossibility trilemma, it has to reduce imports if it aims to lower its national debt by earning trade surpluses and also preserve its competitiveness.
In the following exposition, we first introduce the competitiveness objective in Section 2. Then, Section 3 formulates a necessary condition for a country to launch a trade war or form a trade partnership to optimize its objectives. After that, a bargaining solution is derived in Section 4 to resolve conflicts or share profits. Next, Section 5 discusses a similar condition for globalization or protectionism. Then, in an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, we discuss several extensions and limitations of this framework in Section 7. Our exposition is self-contained and the proofs are in the Appendix.
2 Competitive Advantage in the International Trade System
Based on Adam Smith’s economy of scale and David Ricardo’s comparative advantage, if all production factors were free of flow, then, hopefully, global production would be maximized. However, not all countries share the same well-off status; some succeed over some time, whereas others fail during the same period. Even the most powerful economies have failed in human history, and there needs to be more mechanisms to enforce the fairness of profit sharing in international trade. Thus, a country has to compete for its share and, at the same time, cooperate with others to maximize global production. Both competition and cooperation are essential in international trade.
Domestic production is paramount for policymakers, whereas profit maximization is for firms or producers. Production with zero or slightly negative profit means nothing to firm owners but may still be an option for the government to uplift employment. Production maximization, for example, results in less unemployment, while labor may be cut or resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
We regard the international trade system as a game on a network in which countries act as nodes and trades as directional edges. The trade volumes flow along the edges, and imports and exports move in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or antiglobalization are often framed to shield domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of the trade counterparts.
This section introduces a new type of competitiveness as an objective and social welfare function for the countries in the network. Competitiveness is a relative measure, and we use a constant-sum game to derive it. Absolute production and total trade volume, thus, may give limited information on the economy’s competitive position. The game mocks the in-and-out flows over the network; without data change, these steady flows reach a static equilibrium due to long-run accumulation. The equilibrium is the competitiveness vector, which describes the utilities of the countries’ trade activities. We detail the steps toward the data-based utility and social welfare functions, which are unfamiliar in the trade literature.
2.1 Long-Run Influence in International Trade System
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\not=j$, let $P_{ij}$ be the fraction of country $i$’s production, which exports to country $j$. The exports include all services and goods, such as final goods and intermediate goods, both of which may also contain parts made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are attributable directly to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\not=i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $j\in\mathcal{N}$. We place all the fractions $P_{ij}$ into a $n\times n$ stochastic matrix $P=[P_{ij}]$.
We look for a utility function which determines country $i$’s production or expected production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
$v_{i}(S)\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{j\in S}P_{ij},$ (1)
together with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. So, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. It is normalized since $v_{i}(\mathcal{N})=1$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game and its Shapley value is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$ (Shapley, 1953). Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern (1953) utility function, i.e., $P_{ij}$ is $j$’s utility in the game of $(\mathcal{N},v_{i})$. So, countries have different preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. However, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse judgments of every country in $\mathcal{N}$. It rates the set of countries by $\mathcal{N}$ as a whole.
The element $P_{ij}$ is a type of pro rata power when country $i$ makes its production. The multilinear extension of $v_{i}$ is defined by
$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
where $x_{j}\in[0,1]$ is the probability with which $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also means $j$’s progress in completing its imports from $i$. In addition, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For an annual production, for example, each country finishes $1/12=8.33\%$ of its imports from $i$ in any month. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s influence or power when country $i$ completes its production from 00 to $1$. According to Owen, (1972), the integral equals the Shapley value for the coalitional game, i.e.,
$\int_{0}^{1}\bigtriangledown\tilde{v}_{i}(x,\cdots,x)\mathrm{d}x=(P_{i1},P_{i2},\cdots,P_{in})^{\top}$
where $\top$” denotes vector or matrix transpose.
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them together. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. They are primarily payments for past, current, or future exports. Trade is about exchanging goods and services; as the medium of exchange, monetary flow makes transactions possible in modern international trade. Nevertheless, payments alone could not fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Also, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. Besides, a country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and low-risky investment instruments, so foreign governments or companies must earn extra U.S. dollars from their trades to put into their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for a fictitious example, the chip and design software of the phone could come from the USA and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
in which $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs all countries in $\mathcal{N}$.
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. As a result, $P^{t}$ converges only if $t$ is large enough for all the direct impacts to travel from the top of global economic chains down to the very bottom. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.
2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, called authority distribution in Hu and Shapley, (2003), also solves the following counterbalance equilibrium
$\pi=\pi P$ (3)
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and $\pi_{i}\geq 0$. There exist many counterbalanced systems in the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties — even the most competitive economy needs both cooperation and competition. Between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Though unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). An econometric analysis, for example, could hardly capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From the noncooperative aspect, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase of $\pi_{j}$ may imply a decrease of $\pi_{i}$. So, in theory, there are $n(n-1)/2$ potential trade conflicts, either small or large, in the trade system. Most of them are minor trade disagreements or disputes, so the “war” metaphor is hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Also, the increase of $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initialized by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively.
From the cooperative aspect, country $i$ should help $j$ improve $\pi_{j}$ whenever $P_{ji}>0$ because $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}P_{ji}$ (4)
from Eq. (3). Also, Eq. (3) implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t}$ in which
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}\left(P^{t}\right)_{ji}$ (5)
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. So, Eq. (5) further implies that $i$ should help all other countries in the trade system, including the poorest and the least competitive ones, to enhance their $\pi_{j}$. Thus, the double-edged counterbalance implies that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is how much it should assist or contest $j$ without sacrificing its national interests and even increasing $\pi_{i}$.
[…]
Figure I: Dynamics of authority flow for country $i$ and $j$.
We can physically visualize $\pi_{i}$ as a container or pool in the dynamics of counterbalance equilibrium, in which comparative advantages are the primary driving force for the flows. On the one hand, in the inflow equation Eq. (4), $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, other things being equal. It also derives more authority from countries on which it has a sizable direct influence (i.e., $P_{ji}$ is large), other things being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$ and so, authority flows from $j$ to $i$ alongside the same column. On the other hand, $i$ also contributes to other countries as seen in the following outflow equation
$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\not=i}\pi_{k}P_{kj}.$
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, other things being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries have a break-even point when
$\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ takes more from or gives more to country $j$, respectively. For country $i$, a simple implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\not=j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\not=j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ still maintains an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
2.3 $\pi_{i}$ as Competitive Advantage
Country $i$’s competitive advantage refers to its ability to do better than its competitors in the trade system. For any product exported from country $i$ to $j$, $i$’s competitors include all countries other than itself. So, its competitive advantage extends from bilateral country relations to a relationship with the whole system — it is the ability to beat the best of all other countries in the $j$th market for any $j\in\mathcal{N}$. On the contrary, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over a specific trade partner on particular services or products; it is a bilateral relationship. According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, and a low cost of capital. These factors, imported from the outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker — in Eq. (4), $\pi_{j}P_{ji}$ contains $j$’s return of competitiveness to $\pi_{i}$. In the above smartphone example, the competitive factors include the chips and patents from the USA, the cheap labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means a solid competitiveness to occupy the $j$th market and win a lion share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$ because domestic products also compete with country $j$’s imports when the consumers make their consumption choices, and they have no home bias. We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. The weights are not exogenously assigned but endogenously determined. Also, more weight should be placed on a more competitive reference market than on a less competitive one because it is more challenging to occupy a competitive market than a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. So, both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
Theorem 1
If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
The national competitiveness $\pi_{i}$ is not the sum of the individual competitiveness of the country’s domestic firms, though they are positively associated. In the endogenous weighting, worldwide consumers unconsciously participate in the evaluation of competitiveness, and all products are involved. Essentially, countries unknowingly judge countries themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is, to a great extent, endogenously determined by the distribution of firm productivity in individual countries. Nevertheless, government policies can affect the distribution without significantly mitigating the welfare of international trade by switching imports or exports from one country to another. It can also enhance firm productivity through tax exemption, subsidies, and heavy tariffs on foreign products.
Also, $\pi$ is a von Neumann-Morgenstern Bergson social welfare function (e.g., Harsanyi, 1955; Hammond, 1987) since it is a positively weighted sum of von Neumann-Morgenstern individual utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$; the weights are $\pi$ itself. Thus, the weighted utility function $\pi$ is not determined by how much production factors are consumed but by how much the consumers prefer the products, judged regardless of national interests or preferences by individual countries. Pareto optimality is satisfied when weighing these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant.
However, the welfare function $\pi$ departs significantly from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors but implied from the counterbalance equilibrium because indirect exports also act upon the final products or services. Any exported goods and services from one country to another have different utilities for the exporting and importing countries. Therefore, each country has its unique utility function, and a single utility function or linear regression cannot capture the diversity of all countries. Also, we use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages these factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models. Thus, no unknown parameters are to be estimated, and no residuals exist to differentiate the data from a parametric model. According to Barney and Felin, (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Because of broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network. Also, unjustified assumptions and overly mathematization could mislead the conclusions and policy implications (e.g., Romer, 2015).
The index $\pi_{i}$ contains an inalienable part besides absorbing spillover effects through $P$. There are $t-1$ occurrences of $P_{ii}$ in Eq. (2), which largely accounts for $\pi_{i}P_{ii}$ in Eq. (4) as $t\to\infty$. A global value chain generally takes only finite steps, but it also contains indefinite domestic loops when we trace all the added values of the final products. In the smartphone example, the manufacturing labor’s education, skills, and experience are ultimately achieved domestically over many years, but all somehow contribute to the final product. Also, many USA workers develop the patents and design software domestically for an extended period. These inner loops affect other countries but not vice versa; this part of competitiveness never flows in from external. Thus, social infrastructure and political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that help do business all contribute to national competitiveness (e.g., Delgado et al., 2012). Also, raw oil fields, natural gas reserves, rare minerals, and patented technologies do not come from external sources and remain with $\pi_{i}P_{ii}$ as the nation’s core competitive factors. Refraining exports of rare earth elements or licensed foreign usage of patented technologies, for example, could destroy one or more global value chains.
For a specific global value chain, on the other end of the inalienable parts is the design and assembly of the final product. Product leadership bridges the market, the consumers, strategic direction, technology innovation, sales, and production factors together; this leadership is another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability curtail the risk of costs.
Finally, Eq. (2) implies a solid positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. Nations sell goods that they have in abundance and buy goods that are in short supply domestically. With great abundance and shortage, country $i$ could have vast nonzero elements in the $i$th column and row of the matrix $P$, resulting in astronomical nonzero items summed in Eq. (2). A large economy also produces a high differentiation of goods and services, which bear location advantage and minimal transaction costs for domestic consumers. Free of tariffs and overseas shipping, they also contain a price advantage. For these reasons, $P_{ii}$ is often the largest in the $i$th row of $P$, with or without the home bias. Therefore, a massive economy with a comprehensively integrated industrial system (such as China and the USA) is likely, but not necessary, to have a large value in $\pi$. A country with a single industrial structure is vulnerable to brutal competition, and its $\pi_{i}$ is fragile and volatile when battling with the USA for competitiveness.
Therefore, we use $\pi_{i}$ as an objective for country $i$ when making trade policies. It quantifies the economy’s depth and interplays with breadth; it also incorporates cooperation and non-cooperation with others. It is endogenously extracted from the comparative advantages of all products and services. As each country has a unique location, natural and human resources, and economic systems, a single utility function, which presumes one stereotype for all economies, may be inadequate to capture this diversity. In our study, we recognize that each country, as is its trade stance with partners, is different. Each adds inclusiveness value to the global economy, of which trade between nations is a mainstay.
2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness, i.e.,
$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\not=i}\pi_{j}P_{ji},$ (7)
in which the proportion $P_{ii}$ is generally larger than .7.7.7.7. Thus, $\pi$ has the Matthew effect of accumulated advantage, sometimes summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies have substantial competitiveness compared to less powerful ones, revealing a nonlinear relation between $\pi_{i}$ and $i$’s production $g_{i}$. For example, the USA’s $\pi_{i}$ could be much higher than Japan’s, compared to their production ratio. The amplification also comes from the USA’s relative advantages over other countries, compared to Japan’s. In Eq. (7) with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$,
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\not=\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\not=\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ comes from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, in general, for other countries $j$. Alternatively, the Matthew effect is largely explained by the preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among the countries according to how competitive they already are.
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. So, it has a quadratic growth when $\tilde{g}_{i}$ soars linearly and in a linear regression,
$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i}$ (8)
for some positive $c_{1}$ and $c_{2}$. Therefore, when dealing with a rapidly thriving country $j$, $i$ should mitigate its conflicts with $j$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say,
$g_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i},\quad\mathrm{for}\ \mathrm{some}\ c_{4}>0,$
then $g_{i}$ would be accelerated quadratically, giving rise to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $g_{i}$ would grow exponentially to their maximum capacity. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
Granting extra welfare to competitive economies is another major force that drives trade flows over the network. So, an economic superpower would make all efforts to maintain its superpower status when facing challenges from emerging competitors; maintaining its advantage in $\pi$ avoids the domino effect, the opposite of the Matthew effect. As Copeland (2014) says, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Thus, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. Accordingly, an emerging superpower’s aspiration to assume its rightful place as a superpower would face serious challenges. World War I and II were fiery combats between existing and emerging superpowers, resetting new world orders.
The welfare of competitive advantage leads to an agglomeration effect, which clusters all countries around a handful of economic superpowers. For example, country $i$ considers certain imports from either country $j$ or $k$ where $g_{j}$ is significantly larger than $g_{k}$, other things being equal. If $i$ selects $j$, by Eq. (4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
$\pi_{j}\Delta P_{ji}=\pi_{j}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j}{g_{j}}=\frac{\pi_{j}}{g_{j}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j)$
where $\Delta P_{ji}$ is the change on $P_{ji}$ due to the imports. The approximation ignores the ripple effect from the shock. If $i$ selects $k$, then the competitiveness country $i$ derives from $k$ in the imports is approximately
$\pi_{k}\Delta P_{ki}=\pi_{k}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k}{g_{k}}=\frac{\pi_{k}}{g_{k}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k).$
Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$ and, so, country $i$ would likely choose $j$, other things being equal. Otherwise, $k$ could counter the offer by lowering its price of exports to $(\pi_{k}g_{j})/(\pi_{j}g_{k})$ of $j$’s price. In either case, a small country (such as $k$) tends to fear the emergence of a neighboring superpower (such as $j$) that shares similar location advantage and resource endowment.
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It is a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal such that domestic and foreign goods can compete for the domestic market, and the eigenvector $\pi$ measures the competitiveness. The same domestic products are used domestically or overseas. Hu, (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; so, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Besides, $P$ normalizes the scales of economies such that each row has a unit sum. As a result, $P$ has the largest row eigenvalue $1$, which is common to all stochastic matrices, and its unit-sum eigenvector $\pi$, which is specific to $P$.
3 Bilateral Trade War for Competitiveness
Expecting a positive change to $\pi_{i}$, country $i$ considers a trade war against its trade partner $j$. Gaming on the matrix $P$, $i$’s first-move action could change the element $P_{ji}$ — the portion of $j$’s production exported to $i$. Raising tariffs or decreasing import quotas upon $j$, for example, reduces imports from country $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume that the bilateral conflict involves only the two counterparties without directly sacrificing or benefiting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. Nevertheless, the change on $P_{ji}$ or $P_{ij}$ eventually ramifies over the global value chains in multiple steps, and all third parties get involved indirectly. The simplicity could be far from realistic in the 2018 China-USA trade war as the Biden administration also urged its allies to ban Chinese companies from buying advanced chips and chip-making equipment (e.g., New York Times, 2022).
As a reprisal, country $j$ could react by decreasing $P_{ij}$ or $P_{ji}$ sooner or later. Without the reprisal, the war dies out at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages, but it could also bring in further sanctions from $i$ when $j$ shows a sign of weakness. Overreaction to the hostile change on $P_{ji}$ or $P_{ij}$ may also severely hurt $\pi_{j}$. After the Trump administration began setting tariffs and other trade barriers on China in January 2018, for example, China responded in April 2018 by imposing tariffs on 128 products imported from America (cf., Washington Post, 2018). Say, in retaliation for a $\Delta P_{ji}$ change on $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s $\Delta P_{ij}$ change on $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. So, $\lambda_{ij}=1/\lambda_{ji}$ in the veil of any first-mover advantage. We assume that $\lambda_{ji}>0$ to align the actions from both countries. Thus, ignoring the temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact on $\pi$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; so the reduced exports will be consumed domestically, and the modified $P$ remains as a stochastic matrix. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into a $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$.
The coefficient $\lambda_{ji}$ often depends on many exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. In general, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. For a specific context with a set of fair assumptions, however, many researchers in the literature have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). In the following exposition, we first assume the coefficient is an exogenous constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
We introduce a few notations for the next five theorems and their corollaries. Let $I_{n}$ be the $n\times n$ identity matrix and let $\pi_{{}_{-i}}$ be the transpose of $\pi$ with $\pi_{i}$ removed. The square matrix $Z_{i}$ is the transpose of $P$ with its $i$th row and $i$th column removed. Also, the column vector $\alpha_{i}$ extracts the $i$th row from $P$ and then drops its $i$th element. As usual, $\vec{1}_{n}$ and $\vec{0}_{n}$ are the $n\times 1$ column vectors with all ones and zeros, respectively. For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has one for its $j$th element and zeros elsewhere. Finally, the vector $\gamma_{ji}$ takes the values of $e_{j}$ and then removes its $i$th element. So, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
Theorem 2
With the above setting and notations, for any $j\not=i$,
$\hskip 8.5359pt\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}},$ (9)
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}},$ (10)
and
$\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (11)
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain
$\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
Using Theorem 2, we have the following Corollary 1, already implicated in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
Corollary 1
When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
Besides, since
$(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 2. In particular, when $\lambda_{ji}=0$ in Eq. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Also, Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
Corollary 2
For any $i\not=j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
By varying $j$ in $\mathcal{N}$, we could apply Eq. (9) to identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change on $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$ regardless of the sign of $\mathrm{d}P_{ji}$.
In search for the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
$\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}=\frac{\mathrm{d}\pi_{i}/\pi_{i}}{\mathrm{d}P_{ji}/P_{ji}}=\frac{P_{ji}}{\pi_{i}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}.$ (12)
[…]
$\mathop{\mathrm{argmax}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
as $i$’s best trade partner and
$\mathop{\mathrm{argmin}}\limits_{j\not=i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
as $i$’s best competitor. In Eq. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his best competitor’s best competitor, according to Corollary 2. Thus, $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
Eq. (11) is the side effects on all other countries due to $i$’s change on $P_{ji}$, including $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. With the perturbation on $P_{ji}$, some countries gain, and others lose their competitiveness; the gain or loss may be even more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Though $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not the same magnitude, so one movement could be significant, whereas the other remains insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ unlikely exercises these extreme actions because $\pi_{i}$’s gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been set up to tame trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
4 A Nash Bargaining Solution for Reprisal Coefficients $\Lambda$
When other factors also play a role, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to start a conflict action against $j$. Among them, the most primary are trade deficits. This section studies a bilateral bargaining solution for $\lambda_{ji}$ when trade balance is also an objective. If small economies are willing to dodge direct competition with large ones, the vast majority of the $n(n-1)/2$ conflicts are resolved or mitigated by the bargaining solution.
4.1 Trade Balance
For a given period, country $i$’s trade surplus from $j$ means $j$’s trade deficit of the same magnitude from $i$, and vice versa. With a trade surplus, job opportunities are created in the country, decreasing its unemployment rate and expanding its economy of scale. It also improves the economy’s creditworthiness because it can use the surplus to pay down its foreign debt, if any. A country’s surplus could result from its increasing competitive advantage in tradable goods and services when it beats the competitors in the counterpart’s market. Opinions on the impacts of trade deficits, either bilateral or overall, are diverse and sometimes controversial, although huge deficits can undermine the sustainability of the economy. In the short run, however, trade deficits help nations avoid shortages of goods and mitigate related economic problems, such as inflation and poverty. Implications of trade deficits also depend on their impacts on national security and how they are financed. No nation would completely ignore trade deficits or competitiveness; otherwise, an one-sided emphasis will not last long.
There is a benchmark choice of $\lambda_{ji}$ for net trade balances: $\lambda_{ji}=g_{j}/g_{i}$, the ratio between their productions. In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$ while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$. If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these two countries, as claimed by Theorem 3. This condition is for instant balance change. In aggregate value over a specific time period, there is no trade surplus and deficit between these two countries if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$ because $g_{i}P_{ij}$ is the exports from $i$ to $j$ and $g_{j}P_{ji}$ the exports from $j$ to $i$. For $i$ to have a trade surplus with $j$, we need $g_{i}\lambda_{ji}\Delta P_{ji}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain this trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but hurts $\pi_{i}$ — with overly excessive exports from $i$ to $j$ in the $i$th row of $P$, country $i$ transfers additional production power to $j$, reducing $\pi_{i}$. So $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports even though $i$ loses fewer. This could eventually adversely affect the productions of both countries in the future so that most countries would avoid mutually damaging stalemates. In general, the parameter $\lambda_{ji}$ is determined by both players. However, the sign and size of $\Delta P_{ji}$ are often decided by one only, especially if it has overwhelming power over its counterpart.
Theorem 3
$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
To ease the trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change on $\Delta P_{ji}$. On the other hand, if they look for a balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ removes any trade deficits in the future. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for a long-run rivalry due to $\pi$’s involvement in all countries and all industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and relieve relevant problems such as unemployment and national debt. According to Pozsar, (2022), for example, the 2018 China-USA trade war started in this way: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but built up massive national debt, accumulated through wars on terrors, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China wanted to build a global 5G network and make cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is a proper choice for $\lambda_{ji}$.
We study a general case when a superpower country $i$ faces imminent challenges from an emerging economy $j$. By the Matthew effect, likely $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This can happen only when $\Delta P_{ij}<0$ when both countries lose comparative advantages. Figure II illustrates this trilemma, in which up to two policy objectives are possible for country $i$. If it adopts a position on edge $b$, for example, then it upturns comparative advantage and earns a trade surplus, but gives up competitiveness unless it absorbs competitiveness from other countries. Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned with each other and also do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas $j$’s at another one. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries enjoy escalating economy of scale and diminishing marginal cost of production. According to the proof of Theorem 4, country $j$ could achieve these three objectives listed in the theorem. This seems unfair to country $i$. However, it is conditional on the Matthew effect, which is somewhat “biased” toward $\pi_{i}$. Therefore, Theorem 4 is indeed a redress to the effect.
Theorem 4 (Impossibility Trilemma)
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not achieve all the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
Figure II: Impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
4.2 The Bargaining Solution
We consider a compromise solution between trade balance and competitiveness because a small change $\Delta P_{ji}$ would hurt or benefit both countries. In the trade system, some countries prefer competitiveness while others prefer improved trade balance (either reducing deficit or increasing surplus). Their preferences affect the choice of $\lambda_{ji}$, which holds $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If the countries are indifferent between competitiveness and trade balance, then the middle point seems to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this middle-point solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Otherwise, whenever countries $i$ and $j$ have opposite priorities between enhancing competitiveness and improving trade balance, we have a bargaining problem and apply Nash, (1950) bargaining to find a cooperative solution, which leverages between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ and also binds the condition $\lambda_{ji}\lambda_{ij}=1$.
Let
$p\overset{\mathrm{def}}{=\joinrel=}\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}\quad\mathrm{and}\quad q\overset{\mathrm{def}}{=\joinrel=}\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}$ (15)
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. Without $i$’s bargaining, $\lambda_{ji}$ can secure $p$ and $q$ by Corollary 1 or Theorem 3. Similarly, without $j$’s bargaining, $\lambda_{ij}$ can secure $1/p$ and $1/q$ because
$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}$
and
$\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
We have two scenarios when country $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. So $u_{j}(x)=x$ is an utility function for $j$ because $u_{j}(q)=q>p=u_{j}(p)$. Similarly, $u_{i}(x)=x$ is an utility function for $i$ since $u_{i}(1/p)=1/p>1/q=u_{i}(1/q)$. These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation. By Eq. (15), $p$ and $1/q$ are their respective status quo utilities, obtained if one decides not to bargain with the other player. Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (16)
[…]
Figure III: The Nash bargaining solution $\lambda_{ji}^{*}$ for Eq. (16) and (17), respectively.
In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$. Their utility functions are $u_{j}(x)=q-x$ and $u_{i}(x)=1/p-x$, respectively, and the status quo is 00 for both. The maximization problem becomes
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(q-\lambda_{ji}\right)\left(\frac{1}{p}-\lambda_{ij}\right).$ (17)
The right plot of Figure III illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$ which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$. In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (17) by
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (18)
in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively. To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss. The final outcome is Eq. (18). Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq. (18) and, so, in Eq. (17).
Theorem 5 answers the maximization problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is the same for both preference scenarios, so it also works when country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. In Eq. (16), (17) and (18), both countries have a symmetric role and the same prior bargaining power. This is an axiom of the Nash bargaining. The solution remains unchanged under any positive affine transformation of the utility functions, according to another axiom of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. If $g_{j}<g_{i}$, for example, then $\pi_{j}/\pi_{i}<g_{j}/g_{i}$ is more likely to occur than $\pi_{j}/\pi_{i}\geq g_{j}/g_{i}$, according to the effect. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation fills the gap between these two ratios without changing the likelihood.
Theorem 5
The Nash bargaining solution to Eq. (16), (17), or (18) is
$\left\{\begin{array}[]{rcl}\lambda_{ji}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},\hskip 14.22636pt\mathrm{and}\\ \lambda_{ij}^{*}&\overset{\mathrm{def}}{=\joinrel=}&\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}\end{array}\right.$
for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
The solution of $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$, in general, relies on the choice of status quo points. In Eq. (16), for example, both players could intensify or weaken their status quo positions. If the intensification or weakening is symmetrical with both players, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 3. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
Corollary 3
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$ in Eq. (16), (17), or (18), the solution is still $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
In summary, when country $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{cl}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}&\mathrm{otherwise.}\end{array}\right.$ (19)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb. Country $j$ should avoid competing for a higher $\pi_{j}$ with a significantly larger-sized economy $i$, resulting in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, no matter trade surplus or deficit. This was probably China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$. As a result, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, (1968)’s power transition theory).
Finally, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ in Eq. (19) are also Nash bargaining solutions for $\lambda_{ji}$ because
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (20)
and
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (21)
In Eq. (20) and (21), both countries battle for better competitiveness and net trade balance, respectively. At the solution points, the objective functions are zero and no country gets extra welfare, already implied in Corollary 1 and Theorem 3.
4.3 Posterior Bargaining Power
In a real bargaining context, non-economic factors (such as military power, population, culture, tradition, legislation, and location) and alternative optimal criteria could considerably compromise the symmetry axiom and hence deviate the bargaining solutions substantially different from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$. Thus, the countries likely have unequal prior bargaining power in Theorem 5, and we introduce a posterior bargaining power that updates the prior one.
To extend the bilateral Nash solution in Theorem 5 to multilateral, we define
$\beta_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{\sqrt{\pi_{i}g_{i}}}{\sum\limits_{j\in\mathcal{N}}\sqrt{\pi_{j}g_{j}}}$
as country $i$’s posterior bargaining power or national capability in the trade system. It is normalized such that all bargaining power sums to one and $\beta_{i}$ can be compared over time. Also, it measures bargaining power in multinational negotiations. When allocating votes in forming or reforming a multinational alliance, for example, using the power could avoid any bilateral dialogs and discussions.
Clearly, $\lambda^{*}_{ij}=\beta_{i}/\beta_{j}$ for any $i,j\in\mathcal{N}$. The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ construes a linear ordering for all countries and holds a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of that of $i$ over $j$ and that of $j$ over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. So, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when the companies in country $i$ rely on inputs from its partner countries. The more the companies depend on the partners, the less the country’s power to coerce concessions from the partners. From the consumption aspect, the bargaining power increases when the country consumes more imported goods and services by influencing the production of its counterparts. Between the determinants, $g_{i}$ is the country’s breadth of economy, and $\pi_{i}$ is predominantly the depth of the economy since the matrix $P$ in Eq. (3) already normalizes the production volumes. However, the positive correlation between them implies a risk of domino effect in which a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, which incurs further reduction of $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop of $g_{i}$.
5 Economic Globalization
Globalization refers to removing the barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the United Kingdom’s Brexit in 2020 was a clear rejection of globalization. On the contrary, a protectionism measure could raise tariffs on all imported goods, set import quotas for all other countries, and enact stricter government regulations on all imports. In 2018, for example, the Trump administration imposed punishing tariffs on all steel and aluminum imports from all other countries (e.g., Washington Post, 2018).
Globalization creates winners and losers within a nation but it does not automatically compensate the losers. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. Again, the objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments on the whole matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, say, with a small change $\Delta P_{ii}$ on $P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we make the following assumption: whenever $P$ has a change in the $i$th column, the other columns change proportionally to ensure each row still has a unit sum. Thus, the change $\Delta P_{ii}$ on $P_{ii}$ provokes the adjustments $\Delta P$ on $P$ in three levels of fallout. First, for any $j\not=i$, $P_{ij}$’s proportional response to $\Delta P_{ii}$ is
$\Delta P_{ij}=-\frac{P_{ij}}{1-P_{ii}}\Delta P_{ii},$
and the $i$th row of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{P_{i1}}{1-P_{ii}},\cdots,-\frac{P_{i,i-1}}{1-P_{ii}},1,-\frac{P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{P_{in}}{1-P_{ii}}\right),$
which has a zero sum and the adjustments are proportional to the $i$th row of $P$. Secondly, the change $\Delta P_{ij}$ causes
$\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}=-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}$
due to country $j$’s reprisal reaction. Therefore, the $i$th column of $\Delta P$ is
$\Delta P_{ii}\left(-\frac{\lambda_{i1}P_{i1}}{1-P_{ii}},\cdots,-\frac{\lambda_{i,i-1}P_{i,i-1}}{1-P_{ii}},1,-\frac{\lambda_{i,i+1}P_{i,i+1}}{1-P_{ii}},\cdots,-\frac{\lambda_{in}P_{in}}{1-P_{ii}}\right)^{\top}.$
Lastly, for any $t\not=i$, $P_{jt}$’s proportional response to $\Delta P_{ji}$ is
$\Delta P_{jt}=-\frac{P_{jt}}{1-P_{ji}}\Delta P_{ji}=\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}.$
The sum of the $j$th row in $\Delta P$ is
$-\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\Delta P_{ii}+\sum\limits_{t\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{(1-P_{ii})(1-P_{ji})}\Delta P_{ii}=\frac{\lambda_{ij}P_{ij}}{1-P_{ii}}\left[-1+\sum\limits_{t\not=i}\frac{P_{jt}}{1-P_{ji}}\right]\Delta P_{ii}=0.$
We define the matrix $M$ by
$\left[\begin{array}[]{ccccccc}\frac{\lambda_{i1}P_{i1}P_{11}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1,i-1}}{1-P_{1i}}&-\lambda_{i1}P_{i1}&\frac{\lambda_{i1}P_{i1}P_{1,i+1}}{1-P_{1i}}&\cdots&\frac{\lambda_{i1}P_{i1}P_{1n}}{1-P_{1i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i-1}}{1-P_{i-1,i}}&-\lambda_{i,i-1}P_{i,i-1}&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,i+1}}{1-P_{i-1,i}}&\cdots&\frac{\lambda_{i,i-1}P_{i,i-1}P_{i-1,n}}{1-P_{i-1,i}}\\ -P_{i1}&\cdots&-P_{i,i-1}&1-P_{ii}&-P_{i,i+1}&\cdots&-P_{in}\\ \frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i-1}}{1-P_{i+1,i}}&-\lambda_{i,i+1}P_{i,i+1}&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,i+1}}{1-P_{i+1,i}}&\cdots&\frac{\lambda_{i,i+1}P_{i,i+1}P_{i+1,n}}{1-P_{i+1,i}}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{\lambda_{in}P_{in}P_{n1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{n,i-1}}{1-P_{ni}}&-\lambda_{in}P_{in}&\frac{\lambda_{in}P_{in}P_{n,i+1}}{1-P_{ni}}&\cdots&\frac{\lambda_{in}P_{in}P_{nn}}{1-P_{ni}}\end{array}\right].$
Then $\Delta P=M\Delta P_{ii}/(1-P_{ii})$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M/(1-P_{ii})$ as $\Delta P_{ii}\to 0$. In $M$, there are $22n-$ 2 negative numbers, all in the $i$th row or column; the others are positive. We also let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
Theorem 6
With the above setting and notations, we have
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (22)
and
$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (23)
Thus, the decision rule on globalization could be like this: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and to prevent major internal trade disputes. Also, location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (23) measures the side effects of $i$’s protectionism on the rest of the world. In general, some would benefit from it, whereas others would not.
Country $i$’s optimal level of globalization maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$; the first order condition is necessary at the optimal level. This can simply test if $\pi M_{i}$ is a zero vector or not, according to the first part of Corollary 4. Given $\lambda_{ji}$ for all $i$ and $j$ in $\mathcal{N}$, Eq. (22) and (23) for all $i\in\mathcal{N}$ define a dynamical system that describes the instant change of $\pi$ with respect to the diagonal of $P$. Given the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on $P$. In such an algorithm, for example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% at the same time; and this succession continues until the matrix $P$ converges. At the potential convergence, the final $P$ and $\pi$ reach a new counterbalance equilibrium in Eq. (24). However, $P$ needs no more adjustments if $\Lambda$ satisfies certain desirable properties, such as balanced bilateral trade according to Theorem 3 and the second part of Corollary 4. Also, $P$ needs no adjustments if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or all countries fight for competitiveness in every bilateral conflict. At the convergence, any tiny perturbation on $P_{ii}$ would have a nonsignificant effect on $\pi$, discouraging $i$ from making small changes. Thus, the diagonal of $P$ is relatively stable levels of globalization in these two special cases. Yet, even if a point satisfies all the first-order conditions, it is optimal for some countries only and not every country maximizes its $\pi_{i}$ at this point. Besides, the derivatives in Eq. (22) and (23) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. For these reasons, it might be of limited practical use to seek a far-reached solution to the first-order conditions, even if it exists. Nevertheless, Theorem 6 provides each country with a right direction to steer.
Corollary 4
1. $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$; 2. If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
$\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$ (24)
where $\bigotimes$ is the element-wise matrix product; as two special cases, Eq. (24) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ for all $i$ and $j$ in $\mathcal{N}$, or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (22) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}^{+}$ or $j\in\mathcal{N}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$; and the positive and negative contributions are partially offset. Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as
$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (25)
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is likely better than a pure globalization or protectionism strategy because the right-hand side of Eq. (25) is less than or equal to
$-\sum\limits_{j\in\mathcal{N}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}^{+}$ or $\mathcal{N}^{-}$, respectively.
6 Empirical Study
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$ for the years 2000 to 2019. The bilateral trade data come from the ComTrade database of the United Nations (2021). We use the import and export data not for consumer goods from the World Bank’s WITS Database (2021). Due to the rationality of importers and exporters, the data have largely reflected political, territorial, ideological, cultural, war, national security and other geopolitical considerations, as well as transaction and shipping costs, resource endowment, industrial distribution, and location advantages. For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. During this period, there were a subprime mortgage crisis between 2007 and 2010 and a European debt crisis between 2009 and 2011. All these events heavily impacted relevant countries. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the United Kingdom, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
First, Figure IV plots $\pi_{i}$ from 2000 to 2019. China and Russia gained, and Canada remained almost unchanged while six others lost. Specifically, the USA’s $\pi_{i}$ steadily grew before 2003; it then sharply dropped till 2011 and slowly recovered. Overall, it dropped by 12.7%. China’s $\pi_{i}$ sharply increased during the two crises but almost leveled off elsewhere. Canada and Russia showed their resilience in $\pi$ because of rising prices of commodities in these turmoiling years. The four Western European countries suffered significantly from the debt crisis, and their $\pi_{i}$ continued falling till 2015; manufacturing-focused Germany started falling in 2000 while three others from 2008. Japan’s $\pi_{i}$ plummeted nearly half in the first eight years but finally stopped this momentum between 2007 and 2014. It suffered a domino effect from 2000 to 2007 when China and South Korea challenged its automobile and electric industries. Beyond these nine countries, the others overall became more competitive with their total competitiveness improved by 22.7%, as seen in Figure VI. Many southern asian-, southeast asian-, and middle east countries had tripled their competitiveness including, but not limited to, India, Indonesia, Iraq, United Arab Emirates, and Vietnam.
Also, the Matthew effect was clearly demonstrated. For example, the USA and China’s production ratio was $1.344$ in 2018, but their $\pi_{i}$ ratio was $3.070$. The effect can also be evidenced by Figure VI, in which $\pi_{i}$ has a strong positive correlation with $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
Figure V: The sum of $\pi_{i}$ for all other countries (2000-2019).
Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling $g_{i}$ (2000-2019).
Table I: Country $j$’s direct contribution to $\pi_{i}$ for 2017 and 2019 (in percentage)
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA $\pi_{i}$
CAN $\frac{76.647}{77.390}$ $\frac{1.2635}{1.2732}$ $\frac{0.5390}{0.5179}$ $\frac{0.3222}{0.2781}$ $\frac{0.5922}{0.5728}$ $\frac{0.2877}{0.2607}$ $\frac{0.4329}{0.4768}$ $\frac{0.0410}{0.0355}$ $\frac{14.492}{13.916}$$\frac{.025965}{.026125}$
CHN $\frac{0.3952}{0.3439}$ $\frac{85.584}{84.126}$ $\frac{0.8233}{0.8565}$ $\frac{0.3967}{0.3931}$ $\frac{0.3423}{0.3479}$ $\frac{0.1878}{0.2052}$ $\frac{1.3131}{1.4868}$ $\frac{0.3110}{0.3095}$ $\frac{1.5849}{2.1950}$$\frac{.093905}{.086596}$
DEU $\frac{0.2088}{0.1682}$ $\frac{1.2344}{1.2456}$ $\frac{67.501}{67.410}$ $\frac{2.2469}{2.3993}$ $\frac{1.8879}{2.0095}$ $\frac{1.4928}{1.5634}$ $\frac{0.4265}{0.3507}$ $\frac{0.3737}{0.5379}$ $\frac{3.2504}{3.0282}$$\frac{.040521}{.040105}$
FRA $\frac{0.2345}{0.2171}$ $\frac{0.7032}{0.6617}$ $\frac{3.5649}{3.5687}$ $\frac{74.940}{75.190}$ $\frac{1.9053}{1.8352}$ $\frac{1.6149}{1.5794}$ $\frac{0.2216}{0.2111}$ $\frac{0.1656}{0.2247}$ $\frac{2.4124}{2.3005}$$\frac{.030196}{.030772}$
GBR $\frac{0.6234}{0.6356}$ $\frac{0.9550}{0.9428}$ $\frac{2.2098}{2.4071}$ $\frac{1.5773}{1.6503}$ $\frac{75.873}{75.862}$ $\frac{0.6921}{0.7259}$ $\frac{0.3716}{0.4077}$ $\frac{0.1986}{0.1736}$ $\frac{4.0273}{3.5605}$$\frac{.037385}{.036390}$
ITA $\frac{0.1418}{0.1419}$ $\frac{1.0159}{0.9305}$ $\frac{3.2895}{3.1651}$ $\frac{2.3107}{2.3567}$ $\frac{1.3228}{1.2695}$ $\frac{74.945}{75.662}$ $\frac{0.1958}{0.2123}$ $\frac{0.4274}{0.4525}$ $\frac{1.7055}{1.5402}$$\frac{.019532}{.020507}$
JPN $\frac{0.3297}{0.3118}$ $\frac{2.0275}{1.9912}$ $\frac{0.5242}{0.4869}$ $\frac{0.3184}{0.3041}$ $\frac{0.4209}{0.3614}$ $\frac{0.2218}{0.2139}$ $\frac{87.560}{87.198}$ $\frac{0.1581}{0.1924}$ $\frac{2.2315}{1.9753}$$\frac{.043971}{.045859}$
RUS $\frac{0.1182}{0.1099}$ $\frac{2.6727}{2.2351}$ $\frac{1.7330}{1.7824}$ $\frac{0.9775}{1.0709}$ $\frac{0.7020}{0.7266}$ $\frac{0.8945}{0.8129}$ $\frac{0.6044}{0.5044}$ $\frac{80.225}{79.511}$ $\frac{1.5211}{1.6313}$$\frac{.010657}{.012509}$
USA $\frac{1.3878}{1.4971}$ $\frac{0.8107}{1.0322}$ $\frac{0.4527}{0.4749}$ $\frac{0.2580}{0.2740}$ $\frac{0.5935}{0.5643}$ $\frac{0.1733}{0.1802}$ $\frac{0.3514}{0.3897}$ $\frac{0.0368}{0.0464}$ $\frac{92.142}{91.369}$$\frac{.293880}{.272142}$
* The numerators are for 2019 and the denominators for 2017.
[…]
Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
Figure VII plots the posterior bargaining power from 2000 to 2019. It has similar but smoother patterns as Figure IV. Due to rapid production growth, China’s bargaining power steadily rose, including in 2005 and 2015, when it had troughs in $\pi_{i}$. Also, its $\beta_{i}>\pi_{i}$, whereas the USA’s $\beta_{i}<\pi_{i}$ in each year, showing $\beta_{i}$’s mitigation on the Matthew effect in $\pi_{i}$. Additionally, Russia’s bargaining power sharply dropped after the Russo-Georgian War in 2008 and its annexation of Crimea in 2014. Figure VIII compares the bargaining power versus the USA. The four western European countries’ bargaining power against the USA slightly increased after introducing the Eurozone in 1999 and modestly decreased after 2008. On average, they lost 15.7% of their bargaining power to the USA. Besides, China and Russia improved their relative power multiple times in the first 15 years and remained constant after that. Meanwhile, Japan lost about 50%. Modestly immune to the crises, Canada’s relative bargaining strength sharply rose during the crises and quickly returned to its original level due to its strong alignment with the USA.
Figure VIII: Bargaining power versus the USA (2000-2019).
Table II: $1,000\times\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{-25.78}{-5.78}$ $\frac{-5.49}{-0.37}$ $\frac{-1.47}{0.29}$ $\frac{-0.72}{2.07}$ $\frac{-2.35}{-0.16}$ $\frac{-4.87}{-5.35}$ $\frac{-0.62}{-0.53}$ $\frac{15.75}{63.78}$
CHN $\frac{6.58}{6.90}$ $\frac{7.39}{16.62}$ $\frac{5.36}{7.76}$ $\frac{6.05}{8.46}$ $\frac{2.14}{5.25}$ $\frac{11.56}{31.43}$ $\frac{0.89}{-2.48}$ $\frac{45.19}{65.48}$
DEU $\frac{1.18}{0.14}$ $\frac{-7.73}{-6.05}$ $\frac{6.49}{3.43}$ $\frac{10.07}{5.30}$ $\frac{1.29}{0.18}$ $\frac{0.10}{-5.13}$ $\frac{-1.86}{-4.59}$ $\frac{23.54}{15.06}$
FRA $\frac{0.98}{-0.19}$ $\frac{-8.51}{-5.16}$ $\frac{-13.75}{-6.39}$ $\frac{5.54}{2.67}$ $\frac{-4.20}{-2.57}$ $\frac{-0.90}{-3.67}$ $\frac{-2.08}{-3.10}$ $\frac{12.90}{9.72}$
GBR $\frac{0.73}{-0.96}$ $\frac{-16.47}{-7.66}$ $\frac{-18.11}{-7.01}$ $\frac{-5.31}{-2.28}$ $\frac{-4.64}{-2.58}$ $\frac{-3.39}{-7.65}$ $\frac{-2.38}{-1.99}$ $\frac{8.25}{10.81}$
ITA $\frac{1.01}{0.14}$ $\frac{-8.29}{-5.79}$ $\frac{-3.66}{-0.48}$ $\frac{6.12}{3.66}$ $\frac{7.17}{4.41}$ $\frac{-0.21}{-2.97}$ $\frac{-2.65}{-6.84}$ $\frac{12.51}{11.97}$
JPN $\frac{4.31}{3.52}$ $\frac{-26.28}{-14.31}$ $\frac{-0.28}{4.76}$ $\frac{2.01}{3.72}$ $\frac{4.33}{6.66}$ $\frac{0.35}{2.27}$ $\frac{-1.77}{-2.08}$ $\frac{30.88}{69.74}$
RUS $\frac{1.52}{4.49}$ $\frac{-5.95}{3.48}$ $\frac{8.91}{56.17}$ $\frac{9.85}{24.14}$ $\frac{9.05}{26.41}$ $\frac{5.10}{25.58}$ $\frac{2.78}{6.37}$ $\frac{24.25}{56.37}$
USA $\frac{-1.78}{-5.56}$ $\frac{-48.89}{-17.13}$ $\frac{-12.58}{-6.53}$ $\frac{-4.11}{-2.91}$ $\frac{-3.01}{-2.77}$ $\frac{-4.26}{-2.95}$ $\frac{-10.14}{-23.53}$ $\frac{-2.18}{-1.84}$
* The numerators are for 2017 and the denominators for 2000.
Next, Table II lists the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for 2000 and 2017, calculated from Eq. (9) and (12). If the USA had cut 1% of its imports from China in 2017, for example, its $\pi_{i}$ would have increased by .04889%. While the USA may have had reasons to launch trade wars against all eight others, its top target was China in 2017 and Japan in 2000, according to the numerators and denominators of the last row. As seen from the last column, China and Japan would have been the two largest beneficiaries if they had further collaborated with the USA in 2017 and 2000, respectively. However, comparing the numerators and denominators in the column, the benefits were considerably reduced from 2000 to 2017. Also, starting trade friction from any of these eight countries against the USA would be a loss to the country. Additionally, any other country’s trade war against Russia would be nearly fruitless to increase its $\pi_{i}$, since both numerators and denominators are small in the eighth column. Also, the numbers symmetric with the diagonal have opposite signs, verifying Corollary 2.
Table III: $1,000\times\frac{\mathrm{d}\log\pi_{j}}{\mathrm{d}\log P_{ii}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{52.87}{-188.64}$ $\frac{-16.74}{-9.79}$ $\frac{5.29}{17.08}$ $\frac{4.44}{17.11}$ $\frac{-6.47}{13.46}$ $\frac{5.43}{15.50}$ $\frac{-9.91}{3.11}$ $\frac{3.34}{4.84}$ $\frac{-10.59}{-10.07}$
CHN $\frac{195.55}{35.76}$ $\frac{-2324.27}{-1564.33}$ $\frac{106.47}{31.54}$ $\frac{180.96}{36.26}$ $\frac{232.75}{40.17}$ $\frac{126.87}{31.74}$ $\frac{-81.64}{-33.03}$ $\frac{-135.91}{-2.35}$ $\frac{186.27}{38.81}$
DEU $\frac{38.68}{39.59}$ $\frac{-10.75}{-52.20}$ $\frac{-253.88}{-139.10}$ $\frac{-38.50}{-44.43}$ $\frac{0.50}{-0.04}$ $\frac{-45.19}{-47.50}$ $\frac{14.09}{7.10}$ $\frac{-37.45}{-205.30}$ $\frac{41.38}{44.19}$
FRA $\frac{21.86}{20.97}$ $\frac{-9.83}{-22.47}$ $\frac{-69.45}{-62.18}$ $\frac{-41.92}{-12.66}$ $\frac{-17.07}{-22.86}$ $\frac{-67.71}{-56.16}$ $\frac{1.87}{1.19}$ $\frac{-43.68}{-97.76}$ $\frac{22.93}{25.18}$
GBR $\frac{15.92}{-5.81}$ $\frac{-12.59}{-23.38}$ $\frac{-67.31}{-54.42}$ $\frac{-49.89}{-44.88}$ $\frac{164.52}{145.52}$ $\frac{-42.47}{-29.07}$ $\frac{-7.26}{-13.58}$ $\frac{-33.94}{-117.05}$ $\frac{17.78}{26.54}$
ITA $\frac{18.26}{19.31}$ $\frac{1.06}{-15.52}$ $\frac{-48.34}{-45.02}$ $\frac{-36.50}{-36.91}$ $\frac{-1.16}{-0.96}$ $\frac{-257.39}{-190.00}$ $\frac{4.94}{0.75}$ $\frac{-32.92}{-135.50}$ $\frac{20.83}{23.82}$
JPN $\frac{66.38}{212.93}$ $\frac{-191.11}{-612.51}$ $\frac{35.78}{163.77}$ $\frac{56.09}{212.97}$ $\frac{61.37}{201.14}$ $\frac{46.13}{200.28}$ $\frac{-726.94}{-1518.40}$ $\frac{-19.37}{117.48}$ $\frac{69.31}{236.68}$
RUS $\frac{12.12}{7.85}$ $\frac{-20.29}{-11.83}$ $\frac{-16.42}{4.56}$ $\frac{4.27}{7.57}$ $\frac{8.86}{8.30}$ $\frac{-3.98}{3.36}$ $\frac{-3.65}{3.75}$ $\frac{-832.77}{-2139.12}$ $\frac{12.34}{8.00}$
USA $\frac{-142.03}{-638.28}$ $\frac{-567.12}{-794.20}$ $\frac{-449.08}{-239.99}$ $\frac{-308.90}{-168.36}$ $\frac{-283.70}{-260.03}$ $\frac{-279.33}{-167.92}$ $\frac{-411.84}{-742.34}$ $\frac{-410.17}{-643.02}$ $\frac{834.90}{640.50}$
* The numerators are for 2017 and the denominators are for 2000.
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eq. (22), (23), and (12). From the diagonal, globalization significantly helped China, Germany, Italy, Japan, and Russia. It was a good choice for Canada in 2000 but a bad one in 2017. The effects on France were insignificant because of its small numerator and denominator. The USA and the United Kingdom would be anti-globalization advocates; this explains recent events such as the Trump administration’s withdrawal from international organizations and the United Kingdom’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans and immigrants demanded more inexpensive goods from China, uplifting trade deficits and national debts. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000. The side effects on third parties are off the diagonal. For example, the USA’s further globalization benefited all eight other countries, and so its protectionism would meet serious resistance from its partners. China was the largest beneficiary from the USA’s globalization in both years. Canada suffered the least with the USA’s protectionism in 2017; its further globalization would have slightly profited the USA, according to the first row. The last column implies that further globalization from seven other countries would hurt the USA’s $\pi_{i}$; Japan and China topped the list in 2000 and 2017, respectively. Their hurt declined by 70.2% and ascended by 380%, respectively, in the 17 years. In addition, the second row helps to explain the USA and its western allies’ objection to China’s BRI-alike initiatives, intensified by $4.74$ times from 2000 to 2017 according to the ratios of their numerators and denominators.
Finally, we evaluate the side effects on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is calculated from Eq. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the six other countries were slightly negatively affected in 2018 and 2019. Therefore, these countries would likely take no formal action to prevent the war. Also, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second best would be the United Kingdom for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Probably realizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term battle.
Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
Year CAN DEU FRA GBR ITA JPN RUS CHN USA
2018 -26.114 7.103 3.013 .444 6.661 13.639 15.097 105.808 -53.960
2019 -21.815 3.223 2.133 .039 5.324 12.392 14.894 93.744 -45.954
* Country $i$ is for the USA, $j$ for China, and $k$ for others.
The cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
$\frac{\pi_{j}}{\pi_{i}}=\frac{.092618}{.284327}<\frac{g_{j}}{g_{i}}=\frac{17494.79}{23511.17}.$
For any positive $\mathrm{d}P_{ji}$, therefore, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was
[…]
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. So if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Also, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
7 Discussions
Economic globalization and trade wars have emerged as broad public interests because of global competition among countries, large-scale restructuring of labor markets, and international distribution of added value. We argue that import and export data have already reflected these issues. Therefore, we attempt to extract a country’s competitive strength and bargaining power from the data, identify its best strategies for dealing with trade conflicts and cooperation, and determine if it should move more or less toward globalization. The employed network- and game-theoretical methodologies focus on contemporaneously interactive relationships among economies and are free of parameters and pre-set utility functions. The network game approaches could be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
The forgoing account, however, passes over several vital factors that deserve further investigation. First, counterbalance disequilibrium may exist when trades are notably imbalanced. This could happen in many ways. For example, when country $i$ uses borrowed money to consume $j$’s exports, its consumption still has power over $j$’s production. However, the power comes not only from $i$, but also from the money lender, which has some control over $i$’s consumption. If the lender happens to be $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ underestimated. Secondly, unlike an economic superpower, an emerging economy may focus on its absolute growth of exports instead of competitive advantages. During China’s primitive capital accumulation in the 1980s and 1990s, for example, earning extra US dollars was necessary to finance it industrial modernization. Thus, maximizing $\pi_{i}$ may be a low priority for an emerging economy. The time dimension, absent from the current research, is required to formulate growth. Thirdly, we have yet to include economic globalization’s capital, information, and technology sides. They interplay with goods and services, but their data are not as easily accessible as goods and services. Also, several data issues may cast doubt on the results of the empirical study. There are many missing values in the service data, and country $i$’s reported imports from $j$ are often inconsistent with $j$’s reported exports to $i$. Next, the integration of regional economies merits further attention. Theorem 2 could be used to find the suitable candidates for the integration. If country $i$ belongs to several RTAs, however, $\Delta P_{ii}$ may not be proportionally offset by other countries in Theorem 6. Lastly, one could add other variables to the demand-sided counterbalance Eq. (3), especially those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. Any factors directly impacting $i$’s production could be counted in the $i$th row of $P$. Though consumption may still dominate the fractions, not all productions are driven by consumption.
One could extend the framework in several directions. First, between a bilateral trade war and anti-globalization, a country could target a select group of partners without directly affecting others. For example, as Table II indicates, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan. Secondly, though the dynamical system of Eq. (9)–(11) or Eq. (22)–(23) assumes a fixed $P$, we could include a time dimension to $P$. With the time horizon, we could discuss a trade war’s effects on inflation, $\pi_{i}$’s dependence on its historical values, and the improvement of trade deficits. The time dimension could also improve the forecast capacity of the model. Intertemporal delay often exists when production happens in one year but consumption in the following years. If the production takes two years, then the previous year’s imports of intermediate goods could be included in the current year’s completed production. Next, we could let $\lambda_{ji}$ depend on other determinants besides the competitiveness and production ratios. An industrial or geopolitical analysis could shed light on the choice. For example, as Japan and South Korea rivaled for automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was proper for $\lambda_{ji}$. In the same year, no industry could be a battlefield for China and the USA, so $g_{j}/g_{i}$ seemed a natural choice for $\lambda_{ji}$. Also, in the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954), the distance between trade partners is a vital determinant of their trade flows. Thus, one may discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$. Econometric analyses could also bring additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is larger or less than 00, we may say $\pi_{i}$ is over- or under-valued, respectively. There are other regression forms and $(\pi_{i},g_{i})$ could be panel data. Inclusion of additional regressors, such as debt to GDP ratio, to the regression helps better understand $\pi$ by explanatory data. Finally, anticipating the side effects of $i$’s trade war against $j$, country $j$ could seek cooperation with third parties, which would substantially suffer from the conflict. However, adding more countries to both sides may result in a world war of trade, just like the Cold War between 1947 and 1991.
Reference
Appendix
A1. Proof of Theorem 1
We multiply $\vec{1}_{n}$ on both right sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we then obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$. Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
A2. Proof of Theorem 2
We apply matrix calculus with the restriction of Eq. (3) to prove the theorem. When making a small shock or perturbation $\Delta P$ to $P$ in Eq. (3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation of
$\pi+\Delta\pi=(\pi+\Delta\pi)[P+\Delta P]$ (A.1)
subject to $\Delta P\vec{1}_{n}=\vec{0}_{n}$ and $\Delta\pi\vec{1}_{n}=0$. After subtracting $\pi=\pi P$ from Eq. (A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is
$\Delta\pi[I_{n}-P]\approx\pi\Delta P.$ (A.2)
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{j}$, and $\pi_{-i}$. Accordingly, $P$ have three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the same row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$.
Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\left[I_{n}-P\right]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\left(\vec{0}^{\top}_{i-1},-\lambda_{ji}\pi_{i}+\pi_{j},\vec{0}^{\top}_{j-i-1},\lambda_{ji}\pi_{i}-\pi_{j},\vec{0}^{\top}_{n-j}\right)$ (A.3)
where the right-hand side has all zeros except the $i$th and the $j$th elements.
Also, we partition the transpose of $P$ as
$P^{\top}=\left[\begin{array}[]{ccccc}H_{1}&\eta_{1}&H_{2}&\eta_{2}&H_{3}\\ \mu_{1}&\boxed{P_{ii}}&\mu_{2}&\boxed{P_{ji}}&\mu_{3}\\ H_{4}&\eta_{3}&H_{5}&\eta_{4}&H_{6}\\ \mu_{4}&\boxed{P_{ij}}&\mu_{5}&\boxed{P_{jj}}&\mu_{6}\\ H_{7}&\eta_{5}&H_{8}&\eta_{6}&H_{9}\end{array}\right]$ (A.4)
where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$.
To solve $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ from Eq. (A.3), we write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq. (A.3) as
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&\vec{0}_{i-1}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&-\lambda_{ji}\pi_{i}+\pi_{j}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&\vec{0}_{j-i-1}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\lambda_{ji}\pi_{i}-\pi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&\vec{0}_{n-j}\end{array}\right].$
Since all rows of Eq (A.5), except for the first one, add up to a zero vector, we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{0}&\vec{0}_{j-i-1}^{\top}&\boxed{1}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$ (A.6)
to the left side of Eq. (A.5). Because Eq. (A.6) is the $(n+1)\times(n+1)$ identity matrix except for $n$ ones in the $(i+1)$st row, the multiplication adds the rows of Eq. (A.5) (except for the first row) to its $(i+1)$st row. Consequently, the $(i+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged.
After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$ (A.7)
To make the matrix in Eq. (A.7) lower-triangular, we multiply the following non-singular matrix
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.8)
to the left side of Eq. (A.7) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.9),
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.9),
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
and thus
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}\right]\\ &=&(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
Similarly, after we multiply the following non-singular matrix
$\left[\begin{array}[]{c|c|c|c|c|c}1&\vec{0}_{i-1}^{\top}&0&\vec{0}_{j-i-1}^{\top}&0&\vec{0}_{n-j}^{\top}\\ \vec{0}_{i-1}&I_{i-1}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{j-i-1}^{\top}&\vec{0}_{i-1}&\vec{0}_{i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{0}_{i-1}^{\top}&\boxed{1}&\vec{0}_{j-i-1}^{\top}&\boxed{0}&\vec{0}_{n-j}^{\top}\\ \vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{i-1}^{\top}&\vec{0}_{j-i-1}&I_{j-i-1}&\vec{0}_{j-i-1}&\vec{0}_{j-i-1}\vec{0}_{n-j}^{\top}\\ 0&\vec{1}_{i-1}^{\top}&\boxed{1}&\vec{1}_{j-i-1}^{\top}&\boxed{1}&\vec{1}_{n-j}^{\top}\\ \vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{i-1}^{\top}&\vec{0}_{n-j}&\vec{0}_{n-j}\vec{0}_{j-i-1}^{\top}&\vec{0}_{n-j}&I_{n-j}\end{array}\right]$
to the left side of Eq. (A.5), the $(j+1)$st row of Eq. (A.5) becomes a zero vector and all others remain unchanged. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$ (A.10)
We next multiply the following non-singular matrix
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
to the left side of Eq. (A.10) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}&\vec{0}_{n-1}^{\top}&(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.11),
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}}.$
A3. Proof of Corollary 2
First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices. Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative. Therefore, by Eq. (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$. Also,
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\lambda_{ji}\mathrm{d}P_{ji}}=\frac{1}{\lambda_{ji}^{2}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0.$
A4. Proof of Theorem 4
When $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, there are three possible intervals and two values for $\lambda_{ji}$: $0<\lambda_{ji}<\pi_{j}/\pi_{i}$, $\lambda_{ji}=\pi_{j}/\pi_{i}$, $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$, $\lambda_{ji}=g_{j}/g_{i}$, or $g_{j}/g_{i}<\lambda_{ji}<\infty$. The small change $\Delta P_{ji}$ could be positive or negative, and $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$ has the same sign. Using Theorems 2 and 3, Table V lists all possible outcomes for these ten scenarios of $(\lambda_{ji},\Delta P_{ji})$.
Table V: Scenarios of $(\lambda_{ji},\Delta P_{ji})$ and their impacts on country $i$’s $\pi_{i}$ and net trade with $j$
$0<\lambda_{ji}<\frac{\pi_{j}}{\pi_{i}}$ $\lambda_{ji}=\frac{\pi_{j}}{\pi_{i}}$ $\frac{\pi_{j}}{\pi_{i}}<\lambda_{ji}<\frac{g_{j}}{g_{i}}$ $\lambda_{ji}=\frac{g_{j}}{g_{i}}$ $\frac{g_{j}}{g_{i}}<\lambda_{ji}<\infty$
$\Delta P_{ji}>0$ $\pi_{i}\uparrow$ $\Delta\pi_{i}=0$ $\pi_{i}\downarrow$ $\pi_{i}\downarrow\downarrow$ $\pi_{i}\downarrow\downarrow\downarrow$
---- - - --- - -- zero net $+$
$\Delta P_{ji}<0$ $\pi_{i}\downarrow$ $\Delta\pi_{i}=0$ $\pi_{i}\uparrow$ $\pi_{i}\uparrow\uparrow$ $\pi_{i}\uparrow\uparrow\uparrow$
+++++++ + + +++++ + $+$ zero net --
$+$ and -- for net trade surplus and deficit, respectively.
Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, $+$, or --.
$\Delta\pi_{j}$ and $\Delta\pi_{i}$ have opposite signs and country $i$ and $j$’s net trade balances have opposite signs.
From the first row, country $i$ could not increase $\pi_{i}$ and achieve a trade surplus with country $j$ at the same time. When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$. In these ten scenarios, on the one hand, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes. On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
A5. Proof of Theorem 5
Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (16), the Nash bargaining solution is
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-p\right)\left(\frac{1}{x}-\frac{1}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{q}+\frac{p}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Also by Eq. (17) or (18),
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-q\right)\left(\frac{1}{x}-\frac{1}{p}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{p}+\frac{q}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
A6. Proof of Corollary 3
When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Letting $x=\lambda_{ji}$, the bargaining solution to Eq. (16) becomes
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-cp\right)\left(\frac{1}{x}-\frac{c}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{cx}{q}+\frac{cp}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
In the above, we need $c>0$ and $cp\leq q/c$, i.e., $0<c\leq\sqrt{q/p}$.
After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (17) or (18) becomes
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(\frac{q}{c}-x\right)\left(\frac{1}{cp}-\frac{1}{x}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{q}{cx}+\frac{x}{cp}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Finally, it is easy to check that $\sqrt{pq}\in[cp,q/c]$ for any $0<c\leq\sqrt{q/p}$.
A7. Proof of Theorem 6
With the change $\Delta P_{ii}$ on $P_{ii}$, we have $\Delta P=\frac{\Delta P_{ii}}{1-P_{ii}}M$ and so, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\frac{M}{1-P_{ii}}$ as $\Delta P_{ii}\to 0$. We write $\frac{M}{1-P_{ii}}$ as $\frac{M}{1-P_{ii}}=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $\frac{M}{1-P_{ii}}$. Then $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$. We divide Eq. (A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi\frac{M}{1-P_{ii}}=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.12)
With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.12) is
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&(\pi\xi_{1},\cdots,\pi\xi_{i-1})^{\top}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&\pi\xi_{i}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&(\pi\xi_{i+1},\cdots,\pi\xi_{j-1})^{\top}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\pi\xi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&(\pi\xi_{j+1},\cdots,\pi\xi_{n})^{\top}\end{array}\right].$
As the sum of the last column of (A.13) $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we multiply the nonsingular matrix of Eq. (A.6) to the left side of Eq. (A.13) to make the $(i+1)$st row of Eq. (A.13) to zeros. After dropping the $(i+1)$st row in Eq. (A.13) and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$ as
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$ (A.14)
We next multiply the matrix of Eq. (A.8) to the left side of Eq. (A.14) to get
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\frac{1}{1-P_{ii}}\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}\end{array}\right].$
Therefore, by the first row of Eq. (A.15),
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{1}{1-P_{ii}}\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.15),
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}$
and thus
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[\frac{1}{1-P_{ii}}(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
A8. Proof of Corollary 4
Part 1: If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eq. (22) and Eq. (23). Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
Conversely, if $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$. Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq. (22), we have
$\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}=0.$ (A.16)
We plug Eq. (A.16) into Eq. (23) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get
$\frac{\left(I_{n-1}-Z_{i}\right)^{-1}}{1-P_{ii}}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
Therefore, $\left(\pi M_{i}\right)^{\top}=(1-P_{ii})\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
Part 2: Using the definition of $M_{i}$, $\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that
$\sum\limits_{j\not=i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}P_{it}=0$
for all $t\not=i$. We sum the above equation over all $t\not=i$ to get
[…]
which leads to
$\sum\limits_{j\not=i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.17)
and
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\bigotimes P\right]_{ji}^{\top}=\pi_{i}.$
When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\bigotimes P\right]^{\top}=\pi$.
Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.17) is clearly valid and so is Eq. (24). If $\lambda_{ij}=P_{ji}/P_{ij}$, then Eq. (4) implies Eq. (A.17) and, therefore, Eq. (24) also holds.