### v5

Abstract
This paper explores the exciting dynamics of international trade, focusing on the strategic balance between competition and cooperation in the global trade network. It highlights how competitive advantages, rather than comparative advantages, drive trade conflicts and deglobalization. By using the balance of power, the paper introduces a quantitative measure of competitiveness that serves as a common goal for all countries, alongside trade balance. It then examines how nations can boost their competitive strength and trade balance through globalization, protectionism, trade collaboration, or conflict. The study provides practical insights for policymakers on managing trade relationships, resolving conflicts, and determining the right level of globalization by analyzing trade data. The analysis is supported by a comparison between theory-driven quantitative evidence and historical events, using real-world trade data from 2000 to 2019.
keywords:
globalization, trade war, counterbalance equilibrium, national competitiveness, bargaining power, impossibility trilemma JEL Code: C71 , C78 , F11 , F13 , F15 , O24
Declarations of interests: I have nothing to declare.
Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Declaration of Generative AI and AI-assisted technologies in the writing process: During the preparation of this work, the author never used any Generative AI and AI-assisted technologies.
1 Introduction
Adam Smith’s concept of economies of scale and David Ricardo’s theory of comparative advantage suggest that global production would be maximized if all production factors could move freely. However, not all countries are equally prosperous; some succeed while others struggle. Even powerful economies have failed in the past, highlighting the need for fair profit sharing in international trade. Therefore, countries must compete for their share while cooperating to maximize global production. Both competition and cooperation are essential in international trade. This paper offers fair and strategic solutions for each nation within the trade network.
Analyzing trade wars and globalization presents several challenges. First, international trade involves both conflict and cooperation. While voluntary trade of goods and services results from bilateral collaboration, a country’s growing strength can weaken its trade partner’s position. Thus, a purely cooperative or non-cooperative study may not capture the complexity of trade relationships. The second challenge is creating a high-level description of the interactive trade system that focuses on competitive advantage, as exports and imports already reflect the comparative advantages of specific goods and services. The system is a directional network from which we aim to quantitatively formulate the economies’ utility and social welfare functions in this conflicting and cooperative context. Policy variables, such as imports and exports as percentages of production, should be linked to these economic objectives. Countries can then apply optimization methods to choose appropriate policies to achieve their economic goals. Finally, incorporating data into the analytics can explain past events and find optimal solutions for the near future. The solutions can identify the best trade partner, the right target in trade friction, a fair resolution to a conflict, or the optimal level of globalization.
We view the international trade system as a game on a network where countries act as nodes and trades as directional edges (e.g., Chaney, 2014; $\ddot{\mathrm{O}}$nder and Yilmazkuday, 2016). Trade volumes flow along the edges, with imports and exports moving in opposite directions. In the multilateral network, a unilateral action could achieve unpredictable or even opposite results. Trade wars or anti-globalization efforts are often framed to protect domestic industries or labor markets, reduce trade deficits, or overcome anti-dumping measures. These objectives, however, often result in more harm than gain (e.g., Chung, Lee, Osang, 2016; Read, 2005). Realistically, an attainable goal thrives on mutual understanding of the strengths and weaknesses of trade counterparts. For each country, the strengths and weaknesses lie in its domestic products and foreign imports and exports, forming a coalitional game. The Shapley value (1953) of the game provides a revealed cardinal preference or utility function for the individual country (Roth, 1977) and a power index in completing domestic production.
The common battlefield for competition and cooperation is a social welfare function, which weights the individual utility functions (e.g., Harsanyi, 1955; Hammond, 1987). These weights are chosen to balance the powers within the network (Hu and Shapley, 2003). The social welfare function measures the relative strengths of each country and serves as an objective in our analysis. It integrates both cooperation and competition within the global economic system, assessing the competitive advantages of countries and their long-term impacts on the entire system. This competitiveness generates additional welfare beyond economies of scale. The extra advantage, known as the Matthew effect, is another driving force for collaboration beyond comparative advantages.
De-globalization and trade wars represent two levels of economic conflict. The former involves a country’s withdrawal from the global economy, while the latter involves direct conflict between two trade partners. Both can negatively affect global economic activity. We apply a unified game-theoretic approach to formulate simple yet decisive rules to mitigate either type of conflict. Given that competitiveness has a constant sum in the game, the pursuit of competitiveness becomes the primary reason for initiating a trade war or anti-globalization action, while comparative advantages promote globalization and trade collaboration.
This study focuses on import and export activities, considering that geopolitical factors may already be incorporated into trade behaviors and volumes. Given the rationality of importers and exporters, trade data largely reflect political, territorial, ideological, cultural, war, national security, and other geopolitical considerations, alongside transaction and shipping costs, resource endowment, industrial distribution, and location advantages. From this perspective, a country’s primary objective in engaging in trade wars or de-globalization is to enhance its trade balance and competitiveness, as indicated by the data-driven social welfare function.
The objective of this paper is to offer strategic solutions for policymakers to enhance national competitive strength, rather than focusing solely on the competitiveness of individual firms. Government policymakers, rather than trade companies, are responsible for improving national competitiveness, as trade companies often exploit comparative advantages. A nation is not merely a collection of firms, and the interests of trade companies do not always align with those of policymakers, who must also consider the negative impacts of trade, such as the contraction of domestic labor markets and the relocation of industries overseas. Therefore, our study disregards production-based foundations (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014) and instead maximizes the social welfare function, emphasizing domestic production decision-making by global consumers. However, the measure of competitiveness is also a weighted average of comparative advantages, weighted by the preferences of consumers, whether affluent or impoverished.
When a country initiates a trade war against a trade partner, it raises tariffs or imposes other barriers to enhance competitiveness. However, the outcome depends on the partner’s response. We find that a simple threshold for the counter-reaction — the ratio of their competitiveness — determines whether the country should engage in conflict or cooperation. Additionally, the reaction may benchmark their production ratio to avoid trade deficits. Utilizing the threshold and benchmark ratios, we derive Nash bargaining solutions (1950), generating consistent national bargaining power and providing methods to resolve bilateral trade frictions. A country with no specific targets to confront could choose between further globalization or protectionism. By making minor adjustments to its overall exports, we calibrate the effect on its competitiveness. The sign of the effect indicates whether the country should pursue further globalization. The analyses of trade wars and globalization also shed light on forming multilateral trade agreements.
The solutions are data-driven, relying solely on trade and production data. For a specific context with a set of fair assumptions, many researchers have studied bilateral or even multilateral negotiations (e.g., Bagwell, Staiger, and Yurukoglu, 2020; Schneider, 2005). There are also numerous debates over economic globalization in the literature, with various detailed viewpoints documented in Stiglitz (2002, 2017). Different perspectives, assumptions, and considerations can lead to opposing conclusions about globalization for a country. Our data-driven approach disregards arguments from both sides of the debates about labor markets, production chains, environmental issues, and national security. Trade policymakers may carefully weigh these considerations, and the trade data may already reflect them. Hence, a country could engage in further globalization or protectionism, depending on its particular situation at a specific time.
We test our new theories by analyzing ComTrade data from the United Nations (2021) for the years 2000 through 2019, comparing theory-derived results with major economic events during the same period. Our findings suggest that protectionism could have been a more advantageous strategy for the USA and the United Kingdom (UK), but a less favorable option for China, Germany, Japan, and Russia, particularly if their objectives were centered on competitiveness. Data from 2017 indicate that trade wars might be a viable strategy for the USA to maintain its competitiveness and reduce its trade deficits, and we assess the impacts of the China–USA trade war in 2018 and 2019. However, we also observe significant adverse side effects on third parties. According to our theorem of impossibility trilemma, the USA must reduce imports if it aims to lower its national debt by generating trade surpluses while preserving its competitiveness.
Our contribution to the trade literature lies in introducing a novel strategic and comprehensive framework for analyzing cooperation and competition within the global economic system. Our analyses eschew unnecessary assumptions and excessive mathematization, which can lead to misleading conclusions and policy implications (e.g., Romer, 2015). Unlike traditional studies, our approach does not rely on unknown parameters and formulates a quantitative measure of competitiveness for each country using the balance of power. This measure encompasses overall strength, including both direct and indirect influences. The study also derives national bargaining powers and offers strategic solutions for policymakers to manage trade relationships effectively.
In the following exposition, we first introduce the competitiveness objective in Section 2. Section 3 formulates a necessary condition for a country to initiate a trade war or form a trade partnership to optimize its objectives. Section 4 derives Nash bargaining solutions to resolve bilateral conflicts or share profits. Section 5 discusses a similar condition for globalization or protectionism. In an empirical study of historical events, Section 6 examines the trade data from 2000 to 2019. Finally, Section 7 discusses several extensions and limitations of this framework. Our exposition is self-contained, and the proofs are in the Appendix.
2 A Network-based Social Welfare Function
This section introduces a novel type of competitiveness and social welfare function for countries within the trade network. Competitiveness is a relative measure derived from a constant-sum game. Absolute production and total trade volume may provide limited information regarding an economy’s competitive position. The game simulates the inflows and outflows within the network; without change, these steady flows reach a static equilibrium due to long-term accumulation. The equilibrium represents the competitiveness vector, which describes the utilities of the countries’ positions in the network game. We detail the steps toward the data-based social welfare function and its unique properties, which are unfamiliar in the trade literature.
Domestic production is paramount for policymakers, whereas profit maximization is crucial for firms or producers. Production with zero or slightly negative profit may be insignificant to firm owners but can still be an option for the government to boost employment. For example, production maximization results in less unemployment, while labor may be resourced overseas to maximize shareholders’ profits. Thus, our objective function and strategic responses focus on the production decision-making process, leaving the flow of production factors and the distribution of profit to the firms. Furthermore, the government not only regulates profit-maximization behavior but also improves consumer welfare and redistributes the profits of domestic firms through taxation.
For countries in the network, we describe their aggregate trade activities with a square matrix $P$. Assume there are $n$ countries, labeled as $1,2,\cdots,n$, and denote them collectively by the set $\mathcal{N}=\{1,2,\cdots,n\}$. For any $i,j\in\mathcal{N}$ with $i\neq j$, let $P_{ij}$ be the fraction of country $i$’s production that exports to country $j$. The exports include all services and goods, including intermediate goods; they may contain intermediate goods made by third parties or country $j$. Therefore, the production, denoted by $g_{i}$ for its value, consists of all types of exports as well as all domestically produced final products. Specifically, $g_{i}$ includes country $i$’s GDP and its imports from other countries, which are directly attributable to the production, such as raw materials, intermediate goods, and crude oil. It also includes exports of non-final products but excludes relevant imports, which are already included in $g_{i}$. Besides, $g_{i}$ should exclude exported final goods because they are already included in the GDP. Also, imported final goods which are not directly used in the domestic production should not be counted in $g_{i}$. We use $P_{ii}=1-\sum\limits_{j\neq i}P_{ij}\geq 0$ for the non-exporting fraction. Clearly, all $P_{ij}\geq 0$ for any $i,j\in\mathcal{N}$ and $\sum\limits_{j=1}^{n}P_{ij}=1$ for any $i\in\mathcal{N}$. We place all the fractions $P_{ij}$ into an $n\times n$ stochastic matrix $P=[P_{ij}]$.
Each row of $P$ defines an individual utility function that determines a specific country’s production from a demand- or consumption-side perspective. For any $S\subseteq\mathcal{N}$, we define the set function $v_{i}:2^{\mathcal{N}}\to[0,1]$ by
$v_{i}(S)\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{j\in S}P_{ij},$ (1)
with the convention $v_{i}(\emptyset)=0$ for the empty set $\emptyset$. Thus, $v_{i}(\cdot)$ is a demand-driven production function for country $i$, and $v_{i}(S)$ is its value of production driven by the countries in $S$. Clearly, $(\mathcal{N},v_{i})$ defines a coalitional game, and its Shapley value (1953) is $(P_{i1},P_{i2},\cdots,P_{in})$, the $i$th row of $P$. Roth (1977) shows that the Shapley value is a von Neumann-Morgenstern utility function (1953), i.e., $P_{ij}$ measures an appraisal agent’s satisfaction when they play country $j$’s role in the game of $(\mathcal{N},v_{i})$. The agent is neutral with respect to ordinary and strategic risks detailed in Roth (1977). Thus, countries have cardinal preferences in the game where the inequality $P_{ij}>P_{ik}$ means country $j$ is preferred to $k$. Yet, their preferences change as $i$ varies in $\mathcal{N}$. In the matrix $P$, therefore, each row is an individual utility function, whereas column countries are the arguments or social states of the function. A social welfare function generally combines these individual utility functions into one, accounting for the diverse value judgments of countries in $\mathcal{N}$. It rates the set of countries in $\mathcal{N}$ as a whole and serves as a common platform for the countries to compete and collaborate.
2.1 Long-Run Influence in the Trade System
With the normalization (i.e., $v_{i}(\mathcal{N})=1$ and $v_{i}(\emptyset)=0$) and monotonicity of $v_{i}$, $P_{ij}$ bears much similarity to the Shapley-Shubik power index (1954), measuring country $j$’s chance of pivoting in forming the result $v_{i}(\mathcal{N})$. Indeed, $P_{ij}$ is a type of pro rata power when country $i$ makes its full production. The multilinear extension of $v_{i}$ is a function defined by:
$\tilde{v}_{i}(x_{1},x_{2},\cdots,x_{n})\overset{\mathrm{def}}{=\joinrel=}\sum\limits_{S\subseteq\mathcal{N}}\prod\limits_{j\in S}x_{j}\prod\limits_{k\not\in S}(1-x_{k})v_{i}(S)$
where $x_{j}\in[0,1]$ is the probability with which country $j$ participates in the consumption-driven production. Clearly, when restricted to $\{0,1\}^{n}$ or the corners of $[0,1]^{n}$, $\tilde{v}_{i}$ reduces to $v_{i}$. When the countries act independently, $\tilde{v}_{i}(x_{1},\cdots,x_{n})$ is country $i$’s expected production. Since the production process evolves from zero at the corner $(0,0,\cdots,0)$ to its completion at the corner $(1,1,\cdots,1)$, $x_{j}$ also represents $j$’s progress in completing its imports from $i$. Additionally, the partial derivative of $\tilde{v}_{i}$ with respect to $x_{j}$ measures country $j$’s instant influence or power to increase $\tilde{v}_{i}$ at $(x_{1},x_{2},\cdots,x_{n})$. Without any prior knowledge about the participation probabilities or progresses, we assume they advance evenly along the diagonal from $(0,\cdots,0)$ to $(1,\cdots,1)$. For example, if country $j$ completes 25% of its consumption, all other countries also finish their 25% at the same time. Consequently, the integral of $\tilde{v}_{i}$’s gradient $\bigtriangledown\tilde{v}_{i}$ alongside the diagonal quantifies each country’s overall influence or power that propels country $i$ to complete the production. According to Owen (1972), the integral equals the Shapley value for the coalitional game, i.e.,
$\int_{0}^{1}\bigtriangledown\tilde{v}_{i}(x,\cdots,x)\mathrm{d}x=(P_{i1},P_{i2},\cdots,P_{in})^{\top}$
where $\top$” denotes vector or matrix transpose. Therefore, $P$ contains $n\times n$ quantitative powers, and the balance of power in $P$ refers to a state of stability.
From a supply-side perspective, inputs to the production function in Eq. (1) can include any variables that affect the production and services of country $i$. However, it is impossible to compile a complete list of these variables. Their effects are also time-varying and country-specific; each product has a unique way of combining them. For example, inflows of financial assets, including remittances and foreign direct investments, are influential in making domestic products. These assets are primarily payments for past, current, or future exports. Trade involves exchanging goods and services; and monetary flow as the medium of exchange makes transactions possible in modern international trade. Nevertheless, payments alone cannot fully capture trade activities because exports and imports between two firms may be partially offset, resulting in net payments only. Zero payments may occur for intrafirm transactions of exports within a multinational company. Additionally, trade surpluses and deficits can result from manipulated currency depreciation and appreciation, respectively. A country can accumulate trade deficits if it is a highly desirable destination for foreign investment. For example, U.S. Treasury Bonds are lucrative and risk-free investment instruments after the Cold War, so foreign governments or companies must earn extra U.S. dollars from their trades to add to their foreign investment portfolios.
The direct influences in matrix $P$ automatically generate indirect ones in the closed global trade system. In a chain of production, even though the final product bears the mark of being made in one country, its components or parts may come from elsewhere. When China exports a smartphone to the USA, for example, the chip and design software of the phone could come from the USA, and the camera from Korea. Japan could manufacture the camera screen using US patents. To integrate these spillovers, we consider indirect and long-term impacts by the powers of $P$, such as $P^{2},P^{3},\cdots,P^{\infty}$. In general, $P^{t}$ contains the $t$-step aggregate impacts of all supply chains that crisscross the globalized world. For example, its value at the $i$th row and $j$th column, $\left(P^{t}\right)_{ij}$, accesses all the countries by $t-1$ times because
$\left(P^{t}\right)_{ij}=\sum_{k_{1}=1}^{n}\sum_{k_{2}=1}^{n}\cdots\sum_{k_{t-1}=1}^{n}P_{i,k_{1}}P_{k_{1},k_{2}}P_{k_{2},k_{3}}\cdots P_{k_{t-2},k_{t-1}}P_{k_{t-1},j}$ (2)
where $P_{k_{s},k_{s+1}}$ is the spillover of $k_{s}$’s production to $k_{s+1}$ for any $s=1,2,\cdots,t-1$. It is also $k_{s+1}$’s direct influence on $k_{s}$, which runs through all countries in $\mathcal{N}$.
Under some general conditions (i.e., aperiodicity and irreducibility), the rows of $P^{\infty}$ are a constant row vector $\pi$, which exists uniquely. These conditions are obviously satisfied in the closed and well-connected trade system. Since $\pi_{j}=\lim\limits_{t\to\infty}\left(P^{t}\right)_{ij}$ regardless of the choice of $i$, $\pi_{j}$ measures country $j$’s long-run influence on the whole trade system. The direct influences in $P$ pass through interlinked global value chains and there are millions of them. Consequently, $P^{t}$ converges only if $t$ is large enough. The existence and uniqueness of $\pi$ can be found in the theory of Markov chains (e.g., Karlin and Taylor, 2012). The limit of $P^{t}$ provides a computational method to find $\pi$, and its speed of convergence is linear.
2.2 Mixed Cooperation and Noncooperation
The $1\times n$ row vector $\pi$, referred to as the authority distribution in Hu and Shapley (2003), also solves the following counterbalance equilibrium:
$\pi=\pi P$ (3)
subject to the normalization conditions $\sum\limits_{i=1}^{n}\pi_{i}=1$ and non-negativity condition $\pi_{i}\geq 0$. Numerous counterbalanced systems exist in both the physical and human worlds, such as ecological systems, the USA’s government system, and China’s five-element theory.
The counterbalance in Eq. (3) implies $\pi$’s mixed cooperative and noncooperative properties. This is a distinctive feature between comparative and competitive advantages; while both promote collaborations, competitive advantage may bring trade wars and deglobalization. For instance, between China and the USA from 1980 to 2020, there was more cooperation in the first thirty years than in the last ten years. Although unfamiliar to empirical macroeconomics, this mixture could have an edge in the future (e.g., Allen, 2000, page 147). Econometric analysis often struggles to capture the complexity of antagonism, where any estimated coefficient or implicated effect cannot be both significantly positive and significantly negative.
From a non-cooperative perspective, as $\sum\limits_{i=1}^{n}\pi_{i}=1$, an increase in $\pi_{j}$ may imply a decrease in $\pi_{i}$. Thus, theoretically, there are $n(n-1)/2$ potential trade conflicts, either small or large, within the trade system. Most of these are minor trade disagreements or disputes, making the “war” metaphor hyperbolic in these cases. However, it is worth mitigating even minor disputes before they become problematic or catastrophic. Additionally, the increase in $\pi_{j}$ is not evenly redeemed by other economies; they often wax and wane with some common causality initiated by changes in the matrix $P$. In Sections 3 and 5, we distinguish the first move off the diagonal and that on the diagonal of $P$, respectively, for bilateral trade wars and globalization.
From a cooperative perspective, country $i$ should assist country $j$ in improving $\pi_{j}$ whenever $P_{ji}>0$, as $\pi_{j}P_{ji}$ is part of $\pi_{i}$, as shown in:
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}P_{ji}$ (4)
from Eq. (3). Eq. (3) also implies that $\pi=\pi P=\pi P^{2}=\cdots=\pi P^{t},$ where
$\pi_{i}=\sum\limits_{j=1}^{n}\pi_{j}\left(P^{t}\right)_{ji}$ (5)
for any $t=1,2,3,\cdots.$ When $\pi_{i}>0$, as $\lim\limits_{t\to\infty}\left(P^{t}\right)_{ji}=\pi_{i}>0$, therefore, $\left(P^{t}\right)_{ji}>0$ for all large $t$ and all $j\in\mathcal{N}$. Thus, Eq. (5) further implies that country $i$ should assist all other countries in the trade system, including the poorest and least competitive ones, to enhance their $\pi_{j}$. Hence, the double-edged counterbalance suggests that a carrot-and-stick approach would better leverage country $i$’s position in the trade system when dealing with its trade partner $j$. The trade-off is determining how much it should assist or contest $j$ without sacrificing its national interests and potentially increasing $\pi_{i}$.
[…]
Figure I: Dynamics of authority flow for country $i$ and $j$.
We can conceptualize $\pi_{i}$ as a container or pool within the dynamics of counterbalance equilibrium, where comparative and competitive advantages serve as the primary driving forces for the flows. In the inflow equation Eq. (4), country $i$ accumulates or absorbs authority from other countries through its direct influences. It derives more authority from influential trade partners (i.e., $\pi_{j}$ is large) than from non-influential ones, all else being equal. Additionally, it derives more authority from countries on which it has a significant direct influence (i.e., $P_{ji}$ is large), all else being equal. In Figure I, country $i$ imports goods from $j$ through the $i$th column of $P$, and thus, authority flows from $j$ to $i$ along the same column.
Conversely, country $i$ also contributes to other countries, as illustrated in the following outflow equation:
$\pi_{j}=\pi_{i}P_{ij}+\sum\limits_{k\neq i}\pi_{k}P_{kj}.$
The larger $P_{ij}$ or $\pi_{i}$, the more $i$ contributes to $j$’s authority $\pi_{j}$, all else being equal. In Figure I, country $i$ redistributes $\pi_{i}$ through the $i$th row of $P$, according to its shares of exports.
In the accounting of $\pi$, the two countries reach a break-even point when:
$\pi_{i}P_{ij}=\pi_{j}P_{ji}.$ (6)
Intuitively, if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, then country $i$ either takes more authority from or gives more to country $j$, respectively. For country $i$, a straightforward implication of Eq. (6) is to decrease or increase its imports from $j$ if $\pi_{i}P_{ij}<\pi_{j}P_{ji}$ or $\pi_{i}P_{ij}>\pi_{j}P_{ji}$, respectively. It can also adjust its exports to $j$ accordingly. However, the outflow from $i$ to $j$ does not necessarily match the inflow. With the involvements of third parties, the inflows and outflows meet at the equilibrium Eq. (3) for $i$, i.e. $\pi_{i}P_{ij}+\sum\limits_{k\neq j}\pi_{i}P_{ik}=\pi_{j}P_{ji}+\sum\limits_{k\neq j}\pi_{k}P_{ki}=\pi_{i}$. For the same reason, it could be overly restrictive for all countries to balance their bilateral trades. For example, country $i$ can still maintain an overall zero net balance when it has a trade surplus with $j$ and a trade deficit with $k$ of the same amount.
Finally, Eq. (3) implies that $\pi$ is the unit-sum row eigenvector of the largest row eigenvalue of $P$. It represents a type of eigenvector centrality, a subclass of network centrality. In the literature, eigenvector centrality is the dominant eigenvector for an adjacency or relation matrix, which often has a zero diagonal (e.g., Bonacich, 1987). Our $P$ has a nonzero diagonal, allowing domestic and foreign goods to compete in the domestic market, and the eigenvector $\pi$ measures competitiveness (see Section 2.3). The same domestic products are used both domestically or overseas. Hu (2020) also retains a nonzero diagonal when any accepted student selects only one college to attend from multiple offers; thus, the selected college on the diagonal competes with the others that also accept the student and are off the diagonal. Furthermore, $P$ normalizes the scales of economies such that each row has a unit sum. Consequently, $P$ has the largest row eigenvalue $1$, common to all stochastic matrices.
2.3 $\pi$ as Competitive Advantages
The competitive advantage of country $i$ refers to its ability to outperform its competitors within the trade system. For any product exported from country $i$ to country $j$, the competitors of country $i$ include all other countries. Thus, its competitive advantage extends from bilateral country relations to a relationship with the entire system — it is the ability to surpass all other countries in the $j$th market for any $j\in\mathcal{N}$. In contrast, comparative advantage is an economy’s leverage of specialization, resource endowment, and technology over the trade partner’s opportunity cost; it is a bilateral relationship. Therefore, globalization may have made competitive advantages more prevalent than comparative advantages.
According to the literature (e.g., Porter, 1985; Stutz and Warf, 2010), competitive advantages rest on the direct control of competitive factors of production, such as access to rare resources or advanced technology, cheap but highly skilled labor, land, market affordability, and low cost of capital. These factors, imported from outside or manufactured domestically, are assembled in the final products. Thus, the imported goods carry the original producer’s competitiveness and should be attributed to the original maker. In the smartphone example, the competitive factors include the chips and patents from the USA, the labor from China, the precision manufacturing in Japan, and the sophisticated cameras from Korea. Various studies have also described qualitative strategies to enhance firm competitiveness. For example, Porter (1985) lists three generic strategies: cost leadership, differentiation, and focus.
We find that $\pi$ quantifies national competitiveness. To introduce the quantitative competitiveness, we ignore the home bias in trade (cf., McCallum, 1995) and make a strong assumption for the time being: no bilateral trade deficit nor surplus. Denote $\tau_{j}$ as the $j$th row of $P$. Because of no bilateral surplus nor deficit, $\tau_{j}$ except for its $j$th element equals the imports from other countries to $j$, all in fractions of $j$’s production. So, it measures other countries’ aggregate comparative advantages over country $j$. Besides, all other countries also compete for the comparative advantages over $j$; a large value in $\tau_{j}$ means solid competitiveness to occupy the $j$th market and win a lion’s share of $j$’s imports. If we use the $j$th market as the reference to rank the competitiveness of all countries, then the competitiveness score is the vector $\tau_{j}$, which has a unit sum. Yet, the $j$th value is not excluded from $\tau_{j}$, because domestic products also compete with country $j$’s imports when consumers make their consumption choices, and they have no home bias.
We apply endogenous weighting to the references to combine these scores $\tau_{1},\tau_{2},\cdots,\tau_{n}$. These individual cardinal utilities are not comparable because it is more challenging to occupy a competitive market than a less competitive one. We make them comparable by weighting, and the weights are not exogenously assigned but endogenously determined. More weight should be placed on a more competitive reference market than on a less competitive one. Let the unknown weights be $\rho=(\rho_{1},\rho_{2},\cdots,\rho_{n})$ with $\sum\limits_{j=1}^{n}\rho_{j}=1$ and $\rho_{j}\geq 0$ for all $j\in\mathcal{N}$. Then, we have a weighted competitiveness score $\sum\limits_{j=1}^{n}\rho_{j}\tau_{j}=\rho P$. Both the weighted score and the weight vector quantify the competitiveness of the countries. To be consistent, the weighted average $\rho P$ should be a positive multiple of the weight vector $\rho$. By Theorem 1, $\rho$ is just $\pi$.
Theorem 1
If $\rho P=c\rho$ for some constant $c>0$, then $c=1$ and, therefore, $\rho=\pi$.
In the endogenous weighting, worldwide consumers unconsciously participate in evaluating national competitiveness, encompassing all products and services. Essentially, countries as a whole unknowingly assess themselves without external interventions from business interests, mainstream media, advertising, governments, or even military threats. The structure of the global trade network is largely determined endogenously by the distribution of firm productivity within individual countries. Nevertheless, government policies can influence this distribution without significantly affecting the welfare of international trade by shifting imports or exports between countries. Policies can also enhance firm productivity through tax exemptions, subsidies, and imposing heavy tariffs on foreign products. However, national competitiveness $\pi_{i}$ is not merely the sum of the individual competitiveness of domestic firms; firm competitiveness are weighted in $\pi$ by consumers, wealthy or poor, domestic or overseas. Affordability thus plays a vital role in $\pi$. The positive weights imply a positive correlation between firm and national competitiveness.
Additionally, $\pi$ represents a von Neumann-Morgenstern Bergson social welfare function, as a positively weighted sum of individual von Neumann-Morgenstern utility functions $\tau_{1},\tau_{2},\cdots,\tau_{n}$. Thus, the weighted utility function $\pi$ is determined not by the consumption of production factors or the added value, but by consumer preferences, judged independently of national interests or preferences. Pareto optimality is achieved when weighting these individual utilities: $\pi_{j}\geq\pi_{k}$ if $P_{ij}\geq P_{ik}$ for all $i\in\mathcal{N}$. The optimality remains invariant if any individual utility function is scaled by a positive constant or added by a constant, such as scaling $\tau_{i}$ by $g_{i}$. Harsanyi (1955) and Hammond (1987) provide additional ethical, risk, and utilitarianism connotations for this type of social welfare function. With the weights being $\pi$ itself, $\pi$ exhibits some special properties, discussed in Section 2.4.
However, the social welfare function $\pi$ significantly departs from widely-used utility or welfare functions in the literature on trade wars or globalization (e.g., Grossman and Helpman, 1995; Harrison and Rutstrom, 1991; Ossa, 2014). It is not explicitly expressed in a specific functional form of production factors and total factor productivity but implied from the counterbalance equilibrium, as indirect imports also impact the final products or services. Exported goods and services have different utilities for exporting and importing countries. Therefore, each country has a unique utility function, and a single utility function cannot capture the diversity of all countries. We use fewer assumptions to accomplish $\pi$ by ignoring how efficiently each country packages production factors into final products and how differently it trades the products with others. It is precisely derived from the data and uses no parametric models, thus no unknown parameters need to be estimated, and no residuals exist to differentiate the data from a parametric model. Additionally, according to Barney and Felin (2013), micro-foundations research is still heavily debated, with management, strategy, and organization scholars having varying views on the link between macroeconomic phenomena and microeconomic agents’ behaviors. Due to broad firm and national heterogeneity, micro-founded strategic movements are particularly challenging on the network with flows of millions of products over tens of thousands of directional edges. With this complexity, a certain level of abstraction and ambiguity is necessary to grasp the essential parts of the network.
The index $\pi_{i}$ contains a unique component, $\pi_{i}P_{ii}$, which does not originate from external source, as indicated by Eq. (4). This encompasses, but is not limited to, raw oil fields, natural gas reserves, rare minerals, and patented technologies, all of which remain integral to the nation’s core competitive factors. A global value chain typically involves finite steps but also includes indefinite domestic loops when tracing all the added values of the final products. For instance, in the smartphone example, the education, skills, and experience of manufacturing labor are ultimately developed domestically over many years, contributing to the final product. Additionally, numerous U.S. workers develop patents and design software domestically for extended periods. These inner loops affect other countries but not vice versa; this aspect of competitiveness never flows in from external sources. Consequently, social infrastructure, political institutions, monetary and fiscal policies, microeconomic environments, and other quality determinants that facilitate business all contribute to national competitiveness (e.g., Delgado et al., 2012).
For a specific global value chain, at the other end of the unique component is the design and assembly of the final product. Product leadership bridges the market, consumers, strategic direction, technology innovation, sales, and production factors together; this leadership represents another type of competitiveness. Between the end product and the raw resources in the chain are a series of value-added processes for intermediate goods. Their competitiveness lies in cost efficiency or cost leadership, including the cost of financing. Technological advancements reduce labor costs, and public safety and social stability mitigate the risk of costs.
Eq. (2) indicates a robust positive correlation between $\pi_{i}$ and $g_{i}$, the breadth of the economy. After adjusting for the sizes of $g_{i}$ in $P$, the interaction effects between competitiveness and economies of scale remain, thereby conferring some cost advantages. Nations export goods that are abundant and import goods that are scarce domestically. With significant abundance and shortage, country $i$ could have substantial nonzero elements in the $i$th column and row of the matrix $P$, resulting in considerable nonzero items summed in Eq. (2). A large economy also generates a high differentiation of goods and services, which offer location advantages and minimal transaction costs for domestic consumers. Free from tariffs and overseas shipping, these goods also possess a price advantage. Consequently, $P_{ii}$ is often the largest in the $i$th row of $P$, irrespective of home bias. Therefore, a vast economy with a comprehensively integrated industrial system (such as China and the USA) is likely, though not necessarily, to have a large value in $\pi$. A country with a single industrial structure is susceptible to intense competition, and its $\pi_{i}$ is fragile and volatile when competing with the USA for competitiveness.
Hence, we use $\pi_{i}$ as an objective for country $i$ when formulating trade policies. It quantifies the economy’s depth and its interplay with breadth, encompassing both cooperation and non-cooperation with other nations. This measure is endogenously derived from the comparative advantages of all products and services. Given that each country has unique locations, natural and human resources, and economic systems, a single utility function that assumes one stereotype for all economies may be insufficient to capture this diversity. Our study acknowledges that each country, along with its trade stance with partners, is distinct. Each contributes inclusiveness value to the global economy, where international trade is a cornerstone.
2.4 The Matthew Effect in $\pi$
In Eq. (4), country $i$ accrues its competitiveness in proportion to its initial level of competitiveness:
$\pi_{i}=\pi_{i}P_{ii}+\sum\limits_{j\neq i}\pi_{j}P_{ji},$ (7)
where the proportion $P_{ii}$ is generally larger than .7.7.7.7. This illustrates $\pi$’s Matthew effect of accumulated advantage, often summarized by the aphorism “the rich get richer and the poor get poorer.” Consequently, powerful economies exhibit substantial competitiveness compared to less powerful ones, revealing a quadratic relationship between $\pi_{i}$ and $g_{i}$. For instance, the USA’s $\pi_{i}$ could be significantly higher than Japan’s, relative to their production ratio. The amplification also stems from the USA’s relative advantages over other countries compared to Japan’s. In Eq. (7), with $i=\mathrm{USA}$ or $i=\mathrm{Japan}$, respectively,
$\left\{\begin{array}[]{lcl}\pi_{i}^{\mathrm{USA}}&=&\frac{1}{1-P_{ii}^{\mathrm{USA}}}\sum\limits_{j\neq\mathrm{USA}}\pi_{j}P_{ji}^{\mathrm{USA}},\\ \pi_{i}^{\mathrm{Japan}}&=&\frac{1}{1-P_{ii}^{\mathrm{Japan}}}\sum\limits_{j\neq\mathrm{Japan}}\pi_{j}P_{ji}^{\mathrm{Japan}},\end{array}\right.$
the Mathew effect in $\pi_{i}^{\mathrm{USA}}$ arises from two inequalities: $01>P_{ii}^{\mathrm{USA}}>P_{ii}^{\mathrm{Japan}}>$ 0; $P_{ji}^{\mathrm{USA}}>P_{ji}^{\mathrm{Japan}}>0$, generally for other countries $j$. Alternatively, the Matthew effect is largely explained by preferential attachment in the endogenous weighting, whereby the weight is proportionally distributed among countries according to their existing competitiveness.
The effect implies that $\pi_{i}/\pi_{j}>g_{i}/g_{j}$ when $g_{i}$ is significantly larger than $g_{j}$. If we normalize $g_{i}$ by $\tilde{g}_{i}\overset{\mathrm{def}}{=\joinrel=}g_{i}/\sum\limits_{j=1}^{n}g_{j}$, then the effect also means that $\pi_{i}$ positively correlates with $\tilde{g}_{i}^{2}$ after controlling $\tilde{g}_{i}$. Thus, it exhibits quadratic growth when $\tilde{g}_{i}$ increases linearly, and in a linear regression:
$\pi_{i}=c_{1}\tilde{g}_{i}+c_{2}\tilde{g}_{i}^{2}+\epsilon_{i},$ (8)
for some $c_{1}>0,c_{2}>0$, where $\epsilon_{i}$ is the residual term. Therefore, when dealing with a rapidly thriving country $i$, country $j$ should mitigate its conflicts with $i$ before the counterpart becomes too strong to compromise. Furthermore, if the country capitalizes on the accelerating $\pi_{i}$ in its production $g_{i}$, say:
$\tilde{g}_{i}=c_{3}+c_{4}\pi_{i}+\epsilon_{i}$
for some $c_{4}>0$, then $\tilde{g}_{i}$ would be accelerated quadratically, leading to $\pi_{i}$’s quartic growth. Continuing this iteration, eventually, both $\pi_{i}$ and $\tilde{g}_{i}$ would grow exponentially to their maximum capacities. Many empires have suddenly emerged from the horizon with speedily burgeoning importance in human history.
Additionally, the welfare associated with competitive advantage leads to an agglomeration effect, clustering all countries around a select few of economic superpowers. According to Corollary 1, imports from a large economy are preferred over imports from a smaller economy, all else being equal. Consequently, the smaller economy should counter this preference by lowering its export prices. Meanwhile, a third party could impose different tariffs on large and small countries to ensure that the same imported products have uniform prices in the third country. Furthermore, the large market, characterized by numerous producers and consumers where no single entity can dominate, may be favored by the third party for its market competitiveness. In either scenario, a small country tends to fear the emergence of a neighboring superpower that shares similar location advantages and resource endowments.
Corollary 1
Assume the Matthew effect. If $g_{j}>g_{k}$, then the third party $i$ prefers imports from country $j$ over country $k$, all else being equal.
Thus, providing additional welfare to competitive economies is another significant force driving trade flows across the network, alongside comparative advantage. However, trade wars and protectionism are more likely due to competitive advantages rather than comparative advantages, which benefit all trade partners. Accordingly, an economic superpower would exert all efforts to maintain its status when facing challenges from emerging competitors; preserving its advantage in $\pi$ prevents the domino effect, which is contrary to the Matthew effect. As Copeland (2014) states, “When expectations [of the future trade environment] turn negative, leaders are likely to fear a loss of access to raw materials and markets, giving them an incentive to initiate crises to protect their commercial interests.” Hence, their incentive to start crises is more proportional to $\pi_{i}$ than $g_{i}$. An emerging superpower’s aspiration to assume its rightful place would encounter significant challenges. World War I and II were intense conflicts between existing and emerging superpowers, leading to the establishment of new world orders.
3 Bilateral Trade War for Competitiveness
Anticipating a significant positive change to $\pi_{i}$, country $i$ considers initiating a trade war against its trade partner $j$. Gaming on the matrix $P$, country $i$’s first move could alter the element $P_{ji}$ — the portion of $j$’s production exported to $i$. For instance, raising tariffs or decreasing import quotas on $j$ reduces imports from $j$ and decreases $P_{ji}$. Country $i$ could also restrict certain exports to $j$, directly decreasing $P_{ij}$. For simplicity, we assume the bilateral conflict involves only the two counterparties in $P$ without directly affecting $P_{ik},P_{ki},P_{jk}$, or $P_{kj}$ for any third party $k$. However, changes in $P_{ji}$ or $P_{ij}$ eventually impact global value chains in multiple steps, indirectly involving all third parties. This simplification may not reflect the complexity of the 2018 China-USA trade war, where the Biden administration urged allies to ban Chinese companies from purchasing advanced chips and chip-making equipment (New York Times, 2022).
In retaliation, country $j$ might decrease $P_{ij}$ or $P_{ji}$ sooner or later. Without retaliation, the war ends at inception, and country $i$ wins. No reaction or unconditional surrender could avoid further damages but might also invite additional sanctions from $i$ if $j$ shows weakness. Overreaction to hostile changes in $P_{ji}$ or $P_{ij}$ may also severely harm $\pi_{j}$. For example, after the Trump administration imposed tariffs and trade barriers on China in January 2018, China responded in April 2018 by imposing tariffs on 128 products imported from the USA (Washington Post, 2018).
We introduce a matrix $\Lambda$ for retaliation actions. In retaliation for a change $\Delta P_{ji}$ in $P_{ji}$, country $j$ changes $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$, i.e., $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$. Similarly, in retaliation for $i$’s change $\Delta P_{ij}$ in $P_{ij}$, $j$ changes $P_{ji}$ by $\lambda_{ij}\Delta P_{ij}$, i.e., $\Delta P_{ji}=\lambda_{ij}\Delta P_{ij}$. Thus, $\lambda_{ij}=1/\lambda_{ji}$, in the veil of any first-mover advantage. We assume $\lambda_{ji}>0$ to align actions from both countries. Ignoring temporally lagged reactions in the long-run effects on $\pi$, changing $P_{ji}$ by $\Delta P_{ji}$ and changing $P_{ij}$ by $\lambda_{ji}\Delta P_{ji}$ have the same impact. We let $\lambda_{ii}=1$ and place all $\lambda_{ij}$ into an $n\times n$ matrix $\Lambda=[\lambda_{ij}]$, $\forall i,j\in\mathcal{N}$. Finally, to maintain the unit sum in each row of $P$, we need to deduct $P_{jj}$ by $\Delta P_{ji}$ and $P_{ii}$ by $\Delta P_{ij}$; thus, the reduced exports will be consumed domestically, and the modified $P$ remains a stochastic matrix.
The coefficient $\lambda_{ji}$ often depends on various exogenous factors from both sides, such as political considerations, anti-dumping protection, trade deficits, and specific trade commodities or services. Generally, there is no specific formula for every contingency, and negotiations in practice could involve many rounds of bilateral talks. In the following exposition, we first assume the coefficient is an exogenous or predetermined constant in the potential trade war. Then, Section 4 studies a generic but simple bargaining solution of $\lambda_{ji}$ for most of the $n(n-1)/2$ conflicts. The solution aims to mitigate the conflicts and fairly share the costs rather than eliminate them; it serves the diverse interests of the countries. In the mixed cooperative and noncooperative context, simply maximizing $\lambda_{ji}$ or $\lambda_{ij}$ could be of little interest to both counterparts.
We introduce a few notations for the next five theorems and their corollaries. Let $I_{n}$ be the $n\times n$ identity matrix and let $\pi_{{}_{-i}}$ be the transpose of $\pi$ with $\pi_{i}$ removed. The square matrix $Z_{i}$ is the transpose of $P$ with its $i$th row and $i$th column removed. Also, the column vector $\alpha_{i}$ extracts the $i$th row from $P$ and then drops its $i$th element. As usual, $\vec{1}_{n}$ and $\vec{0}_{n}$ are the $n\times 1$ column vectors with all ones and zeros, respectively. For any $j=1,\cdots,n$, the $n\times 1$ vector $e_{j}$ has one for its $j$th element and zeros elsewhere. Finally, the vector $\gamma_{ji}$ takes the values of $e_{j}$ and then removes its $i$th element. So, $\alpha_{i}$, $\pi_{{}_{-i}}$, and $\gamma_{ji}$ are all non-negative $(n-1)\times 1$ column vectors.
Theorem 2
With the above setting and notations, for any $j\neq i$,
$\hskip 8.5359pt\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}},$ (9)
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}},$ (10)
and
$\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (11)
Capitalizing on the relation $\mathrm{d}P_{ij}=\lambda_{ji}\mathrm{d}P_{ji}$ and $\lambda_{ij}={1}/{\lambda_{ji}}$, we also obtain:
$\left\{\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}&=&\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}=\lambda_{ij}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\\ \frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ij}}&=&\lambda_{ij}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}.\end{array}\right.$
Using Theorem 2, we have the following Corollary 2, already implied in Eq. (6) —  $\pi_{i}$ remains unchanged when changed authority flow from $i$ to $j$ equals that from $j$ to $i$. However, Eq. (6) applies no assumptions in Theorem 2.
Corollary 2
When $\lambda_{ji}=\pi_{j}/\pi_{i}$ or $\lambda_{ij}=\pi_{i}/\pi_{j}$, both $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ij}}$ are a zero vector.
Additionally, since
$(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$
has all non-negative elements, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ and $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$ have opposite signs, as stated in Corollary 3. In particular, when $\lambda_{ji}=0$ in Eqs. (9)-(10), dumping goods and services to country $j$ without resistance would improve $\pi_{i}$ and deteriorate $\pi_{j}$. Eq. (10) may imply that $\lambda_{ji}$ measures $j$’s negotiation power with $i$ in the trade friction if $\pi_{j}$ is $j$’s sole objective.
Corollary 3
For any $i\neq j$, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$ and $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}\leq 0.$
By varying $j$ in Eq. (9), we can identify country $i$’s trade collaborators and competitors. In Eq. (9), $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}$ has an opposite sign of $\lambda_{ji}\pi_{i}-\pi_{j}$. If $\lambda_{ji}>\pi_{j}/\pi_{i}$, then $j$ is a potential target for conflict with $i$ because a negative infinitesimal $\mathrm{d}P_{ji}$ or a negative small $\Delta P_{ji}$ would bring a positive $\mathrm{d}\pi_{i}$ or $\Delta\pi_{i}$, respectively. Similarly, the condition $\lambda_{ji}<\pi_{j}/\pi_{i}$ implies that $j$ is a potential target for further collaboration. Whenever $\lambda_{ji}<\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}>0$, the larger $\lambda_{ji}$, the slower $\pi_{i}$ grows and $\pi_{j}$ shrinks. When $\lambda_{ji}>\pi_{j}/\pi_{i}$, $i$ could also increase $\pi_{i}$ by adjusting the sign of $\mathrm{d}P_{ji}$. If $j$ insists on $\lambda_{ji}=\pi_{j}/\pi_{i}$, then any tiny change in $P_{ji}$ would have no significant effects on both $\pi_{i}$ and $\pi_{j}$, regardless of the sign of $\mathrm{d}P_{ji}$.
In search of the best competitor and collaborator, we make the results comparable across countries by taking percentage changes in both $\pi_{i}$ and $P_{ji}$. The limit of the percentage change of $\pi_{i}$ with respect to that of $P_{ji}$ is
$\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}=\frac{\mathrm{d}\pi_{i}/\pi_{i}}{\mathrm{d}P_{ji}/P_{ji}}=\frac{P_{ji}}{\pi_{i}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}.$ (12)
[…]
Furthermore, country $i$ would choose:
$\mathop{\mathrm{argmax}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (13)
as $i$’s best trade partner and:
$\mathop{\mathrm{argmin}}\limits_{j\neq i}\frac{\mathrm{d}log(\pi_{i})}{\mathrm{d}log(P_{ji})}$ (14)
as $i$’s worst competitor. In Eqs. (13) or (14), $\pi_{i}$ has the largest increase or decrease, respectively, in percentage per one percentage rise of $P_{ji}$. However, $i$ is not his best partner’s best partner nor his worst competitor’s worst competitor, according to Corollary 3. Thus, countries $i$ and $j$ are not in a sports-like competition or zero-sum game unless they are evenly matched. For example, in the 2018 China-USA trade war, both the Trump and Biden administrations actively challenged China, whereas China passively defended itself.
Eq. (11) describes the side effects on all other countries due to country $i$’s change in $P_{ji}$. With the perturbation in $P_{ji}$, some countries gain, while others lose their competitiveness; the gain or loss may be more substantial than $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}$. Although $\pi_{i}$ and $\pi_{j}$ always move in opposite directions, the movements are not of the same magnitude. One movement could be significant, while the other remains insignificant; it is also possible that both movements are insignificant. Third parties unevenly share the difference between these two movements, and their individual or aggregate shares may be significant.
The side effects may play a role in determining $\lambda_{ji}$ when $i$ is much more potent than $j$. The conflict seems resolved if the first-order condition $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{i}$. However, if $\lambda_{ji}=\pi_{j}/\pi_{i}$ maximizes $\pi_{j}$ instead, then in theory, $i$ could extremely counteract and push $\lambda_{ji}$ near zero. In practice, however, $i$ is unlikely to exercise these extreme actions because its gain from $j$ is capped by $\pi_{j}P_{ji}$, as seen from Eq. (4), which could be negligible to $i$. Also, $\pi_{j}P_{ji}$ could be less than the loss of third parties, measured by Eq. (11). So, excessively non-cooperative counteractions could provoke outrage from these countries, and $j$ could ally with them to seek collective bargaining with $i$. It could also pursue an objective other than maximizing $\pi_{j}$, avoiding confrontation with $i$. For example, it can engage in country $i$’s economy by developing complementary industries and focusing on its comparative advantages. For centuries, many norms and standards have been established to tame unilateral trade bullies; today, the World Trade Organization (WTO) deals with trade rules between nations.
4 Nash Bargaining Solutions for Reprisal Coefficients $\Lambda$
When additional factors are considered, the condition $\lambda_{ji}>\pi_{j}/\pi_{i}$ alone is insufficient for country $i$ to initiate a conflict against country $j$. Among these factors, trade deficits are paramount. This section explores bilateral bargaining solutions for $\lambda_{ji}$ when trade balance is also an objective.
4.1 Trade Balance
For a given period, a trade surplus for country $i$ with country $j$ implies a trade deficit of the same magnitude for $j$ with $i$, and vice versa. A trade surplus generates job opportunities, reduces unemployment, and expands the economy of scale. It also enhances creditworthiness by enabling the country to pay down foreign debt. A surplus can result from competitive advantages in tradable goods and services. Opinions on trade deficits vary, though large deficits can undermine economic sustainability. In the short term, trade deficits help avoid shortages and mitigate issues like inflation and poverty. The implications of trade deficits also depend on their impact on national security and how they are financed. No nation can completely ignore trade deficits or competitiveness; a one-sided emphasis will not endure.
There are two benchmark choices for $\lambda_{ji}$ regarding net trade balances, $g_{j}/g_{i}$ and $P_{ij}/P_{ji}$, as stated in Theorem 3. For country $i$ to have a trade surplus with country $j$ due to the change $\Delta P_{ij}$, $g_{i}\Delta P_{ij}>g_{j}\Delta P_{ji}$. Thus, $\lambda_{ji}>g_{j}/g_{i}$ implies $i$’s trade surplus if $\Delta P_{ji}>0$. To sustain trade peace with $\Delta P_{ji}>0$, a large $\lambda_{ji}$ improves $i$’s trade balance but harms $\pi_{i}$. Excessive exports from $i$ to $j$ transfer power to $j$, reducing $\pi_{i}$. Therefore, $\lambda_{ji}$ should be neither too large nor too small. However, $\lambda_{ji}<g_{j}/g_{i}$ also indicates $i$’s surplus if $\Delta P_{ji}<0$. In this case, both $i$ and $j$ lose exports, though $i$ loses fewer. This could adversely affect future production, so most countries avoid mutually damaging stalemates. Generally, $\lambda_{ji}$ is determined by both players, but the sign and size of $\Delta P_{ji}$ are often decided by one, especially if it has overwhelming power.
Theorem 3
$\lambda_{ji}=g_{j}/g_{i}$ or $g_{j}/g_{i}=P_{ij}/P_{ji}$ maintains a zero net trade deficit and surplus between countries $i$ and $j$, at the instant or cumulative level, respectively.
To ease trade tension, we have studied two ways to determine $\lambda_{ji}$. On the one hand, if both sides aim for higher competitiveness, then $\lambda_{ji}=\pi_{j}/\pi_{i}$ is the steady-state solution for any minor change in $\Delta P_{ji}$. On the other hand, if they seek balanced trade, then $\lambda_{ji}=g_{j}/g_{i}$ eliminates any future trade deficits. Generally speaking, $\lambda_{ji}=\pi_{j}/\pi_{i}$ is for long-term rivalry due to $\pi$’s involvement in all countries and industries; $\lambda_{ji}=g_{j}/g_{i}$ could immediately resolve the trade deficit and alleviate related problems such as unemployment and national debt. According to Pozsar (2022), for example, the 2018 China-USA trade war began in this manner: China had grown exponentially over four decades by exporting less expensive goods overseas, whereas the USA lost much of its manufacturing industries but accumulated massive national debt through wars on terror, significant tax cuts, and the 2007 subprime mortgage crisis. Nevertheless, when China sought to build a global 5G network and produce cutting-edge computer chips, the USA attempted to block the move and urged its allies to join the effort. In this example, the focal point is competitiveness rather than trade balance. Therefore, the conflict is expected to last long, and $\pi_{j}/\pi_{i}$ is an appropriate choice for $\lambda_{ji}$.
We examine a scenario where a superpower country $i$ faces imminent challenges from an emerging economy $j$. According to the Matthew effect, it is likely that $\pi_{j}/\pi_{i}<g_{j}/g_{i}$. Based on Theorem 4, country $i$ cannot simultaneously have an increasing $\pi_{i}$ and a net trade surplus while growing its exports to country $j$. This situation occurs only when $\Delta P_{ij}<0$, causing both countries to lose comparative advantages. Figure II illustrates this trilemma, where up to two policy objectives are achievable for country $i$. For instance, if country $i$ adopts a position on edge $b$, it can overturn comparative advantage and earn a trade surplus but sacrifices competitiveness unless it absorbs competitiveness from other countries. In the early 2000s, bilateral exports between China and the USA increased as a percentage of their respective production. According to the trilemma, the USA had to either lose competitiveness, incur trade deficits, or both.
Theorem 4 (Impossibility Trilemma)
Given $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, country $i$ could not simultaneously achieve the three objectives in its bilateral trade with country $j$: increasing $\pi_{i}$; trade surplus; growing exports (i.e., $\Delta P_{ij}>0$).
Figure II: impossibility trilemma of $\pi_{i}\uparrow$, trade surplus, and $\Delta P_{ij}>0$.
Country $i$’s interests in increasing $\pi_{i}$, trade surplus, and $\Delta P_{ij}>0$ are not well aligned and do not completely contradict those of country $j$. In Figure II, country $i$’s optimal policy interests could lie at any point inside the triangle, whereas country $j$’s interests could lie at another point. A common ground is that $\Delta P_{ij}>0$ and $\Delta P_{ji}>0$ when both countries benefit from escalating economies of scale and diminishing marginal costs of production. According to the proof of Theorem 4, country $j$ could simultaneously achieve all three objectives listed in the theorem. This seems unfair to country $i$, but it is conditional on the Matthew effect, which is somewhat biased toward $\pi_{i}$. Therefore, Theorem 4 serves as a redress to the effect.
4.2 Nash Bargaining Solutions
This subsection examines a compromise solution between trade balance and competitiveness. In the trade system, some countries prioritize competitiveness while others aim to improve trade balance, either by reducing deficits or increasing surpluses. Their preferences influence the choice of $\lambda_{ji}$, which maintains $\pi_{i}$ and trade balance unchanged at $\lambda_{ji}=\pi_{j}/\pi_{i}$ and $\lambda_{ji}=g_{j}/g_{i}$, respectively. If countries are indifferent between competitiveness and trade balance, the midpoint appears to be a solution, i.e., $\lambda_{ji}=\left(g_{j}/g_{i}+\pi_{j}/\pi_{i}\right)/2$ and $\lambda_{ij}=\left(g_{i}/g_{j}+\pi_{i}/\pi_{j}\right)/2$. However, this midpoint solution does not satisfy the identity condition $\lambda_{ji}\lambda_{ij}=1$. Alternatively, when countries $i$ and $j$ have opposing priorities between enhancing competitiveness and improving trade balance, the Nash bargaining solution (1950) can be employed to find a cooperative solution that balances $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ while adhering to the condition $\lambda_{ji}\lambda_{ij}=1$.
Let
$p\overset{\mathrm{def}}{=\joinrel=}\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}\quad\mathrm{and}\quad q\overset{\mathrm{def}}{=\joinrel=}\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}$ (15)
which are either $\pi_{j}/\pi_{i}$ or $g_{j}/g_{i}$. By Corollary 2 and Theorem 3, without bargaining, $\lambda_{ji}$ can secure $p$ and $q$, and $\lambda_{ij}$ can secure $1/p$ and $1/q$ because:
$\frac{1}{p}=\frac{1}{\min\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\max\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}\quad\mathrm{and}\quad\frac{1}{q}=\frac{1}{\max\left\{\frac{\pi_{j}}{\pi_{i}},\frac{g_{j}}{g_{i}}\right\}}=\min\left\{\frac{\pi_{i}}{\pi_{j}},\frac{g_{i}}{g_{j}}\right\}.$
We consider two scenarios when countries $i$ and $j$ have opposite preferences on $p$ and $q$. In the first scenario, $j$ prefers $q$ to $p$, whereas $i$ prefers $1/p$ to $1/q$. However, $(q,1/p)$ is not a solution for $(\lambda_{ji},\lambda_{ij})$ since $q\times 1/p\neq 1$. By Eq. (15), $p$ and $1/q$ are their respective status quo points, obtained if one decides not to bargain with the other. In the second scenario of opposite preferences, $j$ prefers $p$ to $q$ while $i$ prefers $1/q$ to $1/p$.
Theorem 5 addresses the bargaining problems, and the solution satisfies the identity $\lambda_{ji}\lambda_{ij}=1$ by design. The solution is consistent across both preference scenarios, thus it remains effective when either country $i$ or $j$ is indifferent between greater competitiveness and improved trade balance. The solution remains unchanged under any positive affine transformation of their utility functions, according to a property of the bargaining solution. By taking the square root of $\pi_{j}/\pi_{i}$, $\lambda_{ji}^{*}$ mitigates the Matthew effect in $\pi$. Since $\lambda_{ji}^{*}$ is between $g_{j}/g_{i}$ and $\pi_{j}/\pi_{i}$, the mitigation bridges the gap between these two ratios.
Theorem 5
When countries $i$ and $j$ have opposite preferences on $p$ and $q$, the Nash bargaining solution is:
$\lambda_{ji}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}\hskip 14.22636pt\mathrm{and}\hskip 14.22636pt\lambda_{ij}^{*}\overset{\mathrm{def}}{=\joinrel=}\sqrt{\frac{\pi_{i}g_{i}}{\pi_{j}g_{j}}}$
for $\lambda_{ji}$ and $\lambda_{ij}$, respectively.
Figure III illustrates the bargaining solution. In the left plot (first scenario), $j$ moves from left to right along the curve $\lambda_{ij}=1/\lambda_{ji}$ while $i$ moves from bottom to top along the same curve. When they meet at the solution point $(\lambda^{*}_{ji},\lambda^{*}_{ij})$, the area of the rectangle cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and the status quo point $(p,1/q)$ is maximized. The right plot (second scenario) illustrates this situation where $j$ moves from $q$ to $p$ and $i$ from $1/p$ to $1/q$ along the curve. They meet at the solution point $(\lambda^{*}_{ji},\lambda_{ij}^{*})$, which maximizes the area of the rectangle, cornered by $(\lambda^{*}_{ji},\lambda^{*}_{ij})$ and $(q,1/p)$.
Figure III: Nash bargaining solution for $\lambda_{ji}^{*}$ when $i$ and $j$ have opposite preferences.
The solutions $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$ rely on the choice of status quo points. The effective bargaining set is $[p,q]$ for $\lambda_{ji}$ and $[1/q,1/p]$ for $\lambda_{ij}$ because country $j$ can secure $p$ (or $q$) and $i$ can secure $1/q$ (or $1/p$) in the first (or second, respectively) scenario. If both players intensify or weaken their status quo positions symmetrically, we scale these two status quo points by a positive number $c\leq\sqrt{q/p}$. Then, $\lambda_{ji}$’s effective bargaining set extends from $[p,q]$ to $[cp,q/c]$ and $\lambda_{ij}$’s extends from $[1/q,1/p]$ to $[c/q,1/(cp)]$. However, the bargaining solution remains unchanged, according to Corollary 4. Letting $c\to 0$, the new effective bargaining sets become $(0,\infty)$ while the symmetry still bounds the solution of $\lambda_{ji}^{*}$ within $[p,q]$ and $\lambda_{ij}^{*}$ within $[1/q,1/p]$.
Corollary 4
For any positive $c\leq\sqrt{q/p}$, if we replace $p$ with $cp$ and $q$ with $q/c$, the bargaining solution remains $\lambda_{ji}=\lambda_{ji}^{*}$ and $\lambda_{ij}=\lambda_{ij}^{*}$.
Additionally, as Corollary 5 states, ${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are also Nash bargaining solutions for $\lambda_{ji}$, as implied in Corollary 2 and Theorem 3. In summary, when countries $i$ and $j$’s preferences are known, we recommend using the following $\lambda_{ji}$:
$\lambda_{ji}=\left\{\begin{array}[]{ll}\frac{\pi_{j}}{\pi_{i}},&\mathrm{if\ both\ prefer\ higher\ competitiveness;}\\ \frac{g_{j}}{g_{i}},&\mathrm{if\ both\ prefer\ better\ trade\ balance;}\\ \sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}},&\mathrm{otherwise.}\end{array}\right.$ (16)
Otherwise, we suggest using $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ as a rule of thumb.
Corollary 5
${\pi_{j}}/{\pi_{i}}$ and ${g_{j}}/{g_{i}}$ are Nash bargaining solutions for $\lambda_{ji}$.
Finally, country $j$ should avoid direct confrontation with a significantly larger-sized economy $i$, according to Corollary 6. This was likely China’s strategy and practice from 1980 to 2010 when dealing with greater economic powers; Deng Xiaoping set its diplomatic policy to “keep a low profile and take no lead” (e.g., Yan, 2014), and trade policy is often used as a tool of foreign policy. Consequently, an even distribution of economic, political, and military capabilities between contending countries is likely to increase the chance of war, and peace is preserved best when there is an imbalance of national capabilities between them (e.g., Organski, 1968).
Corollary 6
Country $j$ should avoid homogeneous preferences with a significantly larger-sized country $i$.
4.3 National Bargaining Power
In deriving the solution $\lambda_{ji}^{*}$, both countries have a symmetric role and identical prior bargaining power, as per the Nash bargaining axiom. However, countries $i$ and $j$ typically possess asymmetric prior bargaining power. For instance, we could maximize $\left(\lambda_{ji}-p\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/q\right)^{1-\zeta_{ij}}$ or $\left(\lambda_{ji}-q\right)^{\zeta_{ij}}\left(\lambda_{ij}-1/p\right)^{1-\zeta_{ij}}$ given some constant $\zeta_{ij}\in(0,1)$ (e.g., Anbarci and Sun, 2013). Without prior knowledge about $\zeta_{ij}$, there is no justification to favor one country over another, resulting in $\lambda^{*}_{ji}$. In a real bargaining context, however, the parameter $\zeta_{ij}$ depends on numerous characteristics between these two, as well as other, countries. Non-economic factors such as military power, population, culture, tradition, legislation, and location, along with alternative optimal criteria, can significantly compromise the symmetry axiom and lead to bargaining solutions that differ substantially from $\lambda_{ji}^{*}$ and $\lambda_{ij}^{*}$.
We introduce a posterior bargaining power that integrates both the symmetric prior and the data, capitalizing on the results in Theorem 5. To extend the bilateral Nash solution in Theorem 5 to multilateral negotiations, we define
$\beta_{i}\overset{\mathrm{def}}{=\joinrel=}\frac{\sqrt{\pi_{i}g_{i}}}{\sum\limits_{j\in\mathcal{N}}\sqrt{\pi_{j}g_{j}}}$
as country $i$’s national bargaining power in the trade system. It is normalized so that all bargaining power sums to one, allowing $\beta_{i}$ to be compared over time. This measure aids multinational negotiations. For instance, when allocating votes in forming or reforming a multinational alliance, using this power can avoid bilateral dialogues and discussions. Despite factors other than production and competitiveness, bilateral trade agreements remain more prevalent than multilateral ones; Yilmazkuday and Yilmazkuday (2014) explored some reasons.
The bargaining power $\{\beta_{i}\}_{i=1}^{n}$ constructs a linear ordering for all countries and maintains a consistent transitivity property for the Nash solution $\lambda_{ij}^{*}$. If $i$ has a negotiation advantage over $j$ and $j$ over $k$, i.e., $\beta_{i}>\beta_{j}$ and $\beta_{j}>\beta_{k}$, then $i$ necessarily has more bargaining power than $k$ since $\beta_{i}>\beta_{k}$. Specifically, $i$’s bargaining power over $k$ equals the ratio of $i$’s power over $j$ and $j$’s power over $k$, i.e., $\lambda_{ik}^{*}=\lambda_{ij}^{*}/\lambda_{jk}^{*}$ for any $j\in\mathcal{N}$. Also, we can recover it directly from the relation $\lambda^{*}_{ik}=\beta_{i}/\beta_{k}$.
Furthermore, there are two equally weighted and highly correlated determinants in $\beta_{i}$: the volume of production $g_{i}$ and the strength of competitiveness $\pi_{i}$. Thus, $\beta_{i}$ inherits some “controversial” properties from $\pi_{i}$. From the production aspect, $\beta_{i}$ declines when companies in country $i$ rely on inputs from partner countries. The more companies depend on partners, the less power the country has to coerce concessions from them. From the consumption aspect, bargaining power increases when the country consumes more imported goods and services, thereby influencing the production of its counterparts. Between the determinants, $g_{i}$ represents the country’s breadth of economy, while $\pi_{i}$ predominantly reflects the depth of the economy, as the matrix $P$ in Eq. (3) already normalizes production volumes. However, the positive correlation between them implies a risk of a domino effect, where a decline of $g_{i}$ triggers a decrease of $\pi_{i}$, leading to further reduction in $g_{i}$. Consequently, $\beta_{i}$ would shrink more than the initial drop in $g_{i}$.
5 Economic Globalization
Globalization refers to the removal of barriers to the flow of financial products, goods, technology, information, and jobs across national borders and cultures. In the economic context, it describes the increasing interdependence of a country with the rest of the world, fostered through the trade of goods and services. A recent example is China’s entry into the WTO in 2001, while the UK’s Brexit in 2020 was a clear rejection of globalization. Conversely, protectionist measures could raise tariffs on imported goods, set import quotas for other countries, and enact stricter government regulations on imports. For instance, in 2018, the Trump administration imposed punitive tariffs on all steel and aluminum imports from other countries (Washington Post, 2018).
Globalization creates winners and losers within a nation but does not automatically compensate the losers, thereby fueling income inequality and political divisions. Therefore, a more inclusive globalization policy is needed, one that not only maximizes the winners’ profits but also considers the negative impacts on the losers. This section questions whether the country should retreat from global economic integration without directly attacking any specific country. The objective is to maximize the country’s national competitiveness rather than the aggregate utility function of its trade firms. We consider actions that directly affect the rest of the world and neglect bilateral trade balances.
Country $i$’s globalization action triggers adjustments in the entire matrix $P$. Gaming on $P$, the country’s first move is on the diagonal element $P_{ii}$, with a small change $\Delta P_{ii}$. If it increases globalization, it shrinks its value of $P_{ii}$ and $\Delta P_{ii}<0$; otherwise, it increases $P_{ii}$ and $\Delta P_{ii}>0$ if it boosts protectionism. As the action has no specific target countries, we assume that whenever $P$ changes the $i$th column, the other columns change proportionally to ensure each row of $P$ still has a unit sum. For example, the change $\Delta P_{ji}$ causes proportional changes $\Delta P_{ji}\sigma_{ji}$ to the $j$th row where:
$\sigma_{ji}\overset{\mathrm{def}}{=\joinrel=}\frac{1}{1-P_{ji}}\left(-P_{j1},\cdots,-P_{j,i-1},1-P_{ji},-P_{j,i+1},\cdots,-P_{jn}\right).$
Consequently, the change $\Delta P_{ii}$ provokes adjustments $\Delta P$ in three levels of fallout. First, the $i$th row in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}$. Secondly, the $i$th column in $\Delta P$ is $\Delta P_{ii}\sigma_{ii}\odot\Lambda_{i}$, due to reprisal reactions, where $\odot$ denotes the element-wise product or Hadamard product and $\Lambda_{i}$ is the $i$th row of $\Lambda$. Lastly, the $i$th column of $\Delta P$ triggers changes in all other columns. We define:
$M\overset{\mathrm{def}}{=\joinrel=}\mathrm{diag}\left(\sigma_{ii}\odot\Lambda_{i}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right]$
where $\mathrm{diag}(\sigma_{ii}\odot\Lambda_{i})$ is the diagonal matrix with the diagonal vector $\sigma_{ii}\odot\Lambda_{i}$. Thus, the $j$th row of $M$ is $\sigma_{ji}$, scaled by the $j$th element of $\sigma_{ii}\odot\Lambda_{i}$, and thereby $\Delta P=\Delta P_{ii}M$. So, $\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=M$ as $\Delta P_{ii}\to 0$. We let $M_{i}$ be $M$ with its $i$th column dropped. Theorem 6 lists the derivative of $\pi$ with respect to $P_{ii}$.
Theorem 6
With the above setting and notations, we have:
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}$ (17)
and
$\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\left(I_{n-1}-Z_{i}\right)^{-1}\left[\left(\pi M_{i}\right)^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].$ (18)
The optimal level of globalization for country $i$ maximizes $\pi_{i}$ by adjusting $P_{ii}$ and, accordingly, rebalancing $P$. The first-order condition $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ is necessary at the optimal level because $P_{ii}$ cannot be $1$ or 00. However, $\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ii}}$ may be positive or negative at this optimal level, allowing country $j$ to capitalize on this by making a corresponding negative or positive change in $P_{ij}$, thereby affecting $P_{ii}$. To disincentive other countries from making simultaneous movements, we set $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}$ to be a zero vector. This can be tested by checking if $\pi M_{i}$ is a zero vector, according to Corollary 7.
Corollary 7
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ if and only if $\pi M_{i}=\vec{0}^{\top}_{n-1}$.
Given $\Lambda$, Eqs. (17) and (18) for all $i\in\mathcal{N}$ define a dynamical system describing the instantaneous change of $\pi$ with respect to the diagonal of $P$. There are $n$ independent variables in the system since the proportions among off-diagonal elements remain unchanged. Due to the system’s complexity, numerical algorithms may achieve a solution to $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$ for all $i\in\mathcal{N}$ through successive tiny adjustments on the diagonal. For example, over-globalized countries increase $P_{ii}$ by 1% at each adjustment while under-globalized countries reduce $P_{ii}$ by 1% simultaneously; this succession continues until the matrix $P$ converges.
Alternatively, we study the properties of the solution to the first-order conditions. If the solution exists, it satisfies a new counterbalance equilibrium in Eq. (19), as stated in Corollary 8. Eq. (19) places $n-1$ identities to the solution, requiring only one more condition to uniquely identify it. Also, $P$ and $\pi$ already satisfy the new counterbalance if $\Lambda$ exhibits some desirable properties, such as $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $P_{ji}/P_{ij}$, resulting in bilateral authority or trade balance, according to Eq. (6) or Theorem 3, respectively. At the stable solution, any tiny perturbation on the diagonal of $P$ would have a nonsignificant effect on $\pi$, discouraging countries from making small movement. Yet, this solution is optimal for some countries only, and not every country maximizes its competitiveness with this solution. Additionally, the derivatives in Eqs. (17) and (18) are based on the exogeneity assumption of $\Lambda$ and the proportional rebalancing in $M$. Therefore, seeking a far-reaching solution to the first-order conditions might be of limited practical use. Nevertheless, Theorem 6 provides each country with the right direction to steer.
Corollary 8
If $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}^{\top}_{n}$ for all $i\in\mathcal{N}$, then $\pi$ satisfies that
$\pi\left[\Lambda\odot P\right]^{\top}=\pi.$ (19)
Eq. (19) is satisfied if $\lambda_{ij}=\pi_{i}/\pi_{j}$ or $\lambda_{ij}=P_{ji}/P_{ij}$ for all $i$ and $j$ in $\mathcal{N}$.
The decision rule on globalization could be as follows: country $i$ would engage in further globalization or protectionism if $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is significantly negative or positive, respectively. Otherwise, when $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$ is insignificantly different from zero, country $i$ could seek targets of trade wars or trade partnerships as described in Sections 3 and 4. In between, it could form or participate in preferential trade agreements (PTAs) or regional trade agreements (RTA) with other economies. Within an ideal PTA or RTA, the ebb and flow of $\lambda_{j}/\lambda_{i}$ and $g_{j}/g_{i}$ should be well balanced to facilitate common agreements and prevent major internal trade disputes. Location and cultural proximity often foster the formation of a PTA or RTA. Besides, Eq. (18) measures the side effects of $i$’s protectionism on the rest of the world, helping select countries for a PTA or RTA. In general, some would benefit from it, whereas others would not.
Finally, Theorem 6 implies a mixed globalization and protectionism strategy, helping $i$ to decide which PTA or RTA to join. Let us divide the countries in $\mathcal{N}\setminus\{i\}$ into two groups based on their signs in the vector $\pi M_{i}$: country $j$ belongs to the group of $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$ if its component in $\pi M_{i}$ is positive or negative, respectively. Because $(I_{n-1}-Z_{i})^{-1}$ is a non-negative matrix and $\alpha_{i}$ a non-negative vector, we rewrite Eq. (17) as $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-(\pi M_{i})\omega$ for some non-negative vector $\omega$. Thus, when $j\in\mathcal{N}_{i}^{+}$ or $j\in\mathcal{N}_{i}^{-}$, it contributes negatively or positively, respectively, to $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}$, and the positive and negative contributions are partially offset.
Given $\Delta P_{ii}$, we have the approximate change of $\pi_{i}$ as:
$\Delta\pi_{i}\approx-\sum_{j\in\mathcal{N}_{i}^{+}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}-\sum_{j\in\mathcal{N}_{i}^{-}}\left(\pi M_{i}\right)_{{}_{j}}\omega_{j}\Delta P_{ii}.$ (20)
In the mixed strategy, country $i$ engages in further protectionism with countries in $\mathcal{N}_{i}^{-}$ and further globalization with countries in $\mathcal{N}_{i}^{+}$, avoiding the partial offsetting. This strategy is better than a pure globalization or protectionism strategy because the right-hand side of Eq. (20) is less than or equal to:
$-\sum\limits_{j\in\mathcal{N}_{i}^{+}}(\pi M_{i})_{{}_{j}}\omega_{j}(-|\Delta P_{ii}|)-\sum\limits_{j\in\mathcal{N}_{i}^{-}}(\pi M_{i})_{{}_{j}}\omega_{j}|\Delta P_{ii}|.$
Therefore, country $i$ should increase or decrease cooperation with countries in $\mathcal{N}_{i}^{+}$ or $\mathcal{N}_{i}^{-}$, respectively. If most country members of a PTA or RTA are in $\mathcal{N}_{i}^{+}$, then country $i$ should consider participating in the trade agreement. However, country $j\in\mathcal{N}_{i}^{+}$ does not necessarily imply that country $i\in\mathcal{N}_{j}^{+}$, possibly causing additional negotiations if $j$ is already in the PTA or RTA.
6 Empirical Study
In this empirical study, we apply our theoretical results to real-world trade data spanning from 2000 to 2019. This analysis provides quantitative evidence to support our previous arguments, making the theoretical insights more tangible and relevant for policymakers. The empirical analysis reveals trends in national competitiveness and bargaining power over time, highlighting how different countries have gained or lost. This provides a historical perspective on the impacts of globalization and trade wars.
Between 2000 and 2019, the world economy underwent significant transformations and challenges, albeit with varying degrees of success across different regions. This period witnessed a substantial increase in global trade volumes, driven by rapid technological advancements, the expansion of international trade agreements, and the rise of emerging markets. The integration of China into the WTO in 2001 was a pivotal moment, leading to a surge in global trade activities. The rise of emerging markets, particularly China and India, played a crucial role in global economic growth but also led to a trade war between China and the USA. The 2008 financial crisis and the subsequent European debt crisis were major turning points, causing widespread economic downturns and leading to substantial policy interventions by governments and central banks worldwide.
As trade volumes increased with globalization, nations demanded more of the world’s reserve currency, primarily through trade surpluses. Consequently, the reserve currency appreciates, decreasing the currency issuer’s exports and increasing trade deficits. The issuer may have three options: encouraging local currency settlement in international transactions or reducing its currency’s usage as the world’s reserve currency; ignoring the trade deficits and focusing on national competitiveness while still increasing exports; or rolling back globalization worldwide. For the USA, reshoring its manufacturing sectors would enhance its competitiveness and grow exports; however, trade deficit may still be an issue according to the impossibility trilemma. For China, significant growth is necessary to shrink the competitiveness gap with the USA, according to the Matthew effect in $\pi$.
This empirical analysis calculates $\pi$ and its derivatives with respect to $P_{ii}$ and $P_{ji}$. The bilateral trade data are sourced from the United Nations ComTrade database (2021). We use import and export data for non-consumer goods from the World Bank’s WITS Database (2021). For comparison, we focus on the years 2000, before China entered the WTO in 2001, and 2017, before the beginning of the China-USA trade war in 2018. Though all economies in the data are used in the calculations, we only present the results for China, Russia, and the G7 countries (Canada, France, Germany, Italy, Japan, the UK, and the USA). Their ISO-3 country codes are CHN, RUS, CAN, FRA, DEU, ITA, JPN, GBR, and USA, respectively. We use $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$ in the computations. The results are in the following tables and figures.
Figure IV: Competitiveness $\pi_{i}$ for China, Russia, and G7 countries (2000-2019).
Figure IV illustrates the competitiveness trends from 2000 to 2019. During this period, China and Russia experienced significant gains, while Canada remained relatively stable, and the other six countries saw declines. Notably, the USA’s competitiveness grew steadily until 2003, followed by a sharp decline until 2011, and a gradual recovery thereafter, resulting in an overall decrease of 12.7%. China’s competitiveness surged during the two crises but plateaued otherwise. Canada and Russia demonstrated resilience due to rising commodity prices during transformative years. The four Western European countries were significantly impacted by the debt crisis, with their competitiveness continuing to fall until 2015. Germany, focused on manufacturing, began its decline in 2000, while the other three countries started in 2008. Japan’s competitiveness nearly halved in the first eight years but stabilized between 2007 and 2014, suffering a domino effect from 2000 to 2007 as China and South Korea challenged its automobile and electric industries. Beyond these nine countries, others collectively improved their competitiveness by 22.7%, as depicted in Figure VI. Many Southern Asian, Southeast Asian, and Middle East countries, including India, Indonesia, Iraq, United Arab Emirates, and Vietnam, tripled their competitiveness.
The Matthew effect is clearly demonstrated. For instance, the USA and China’s production ratio was $1.344$ in 2018, but their competitiveness ratio was $3.070$. This effect is further evidenced by Figure VI, which shows a strong positive correlation between $\pi_{i}$ and $g_{i}^{2}$ after controlling $g_{i}$. The partial correlation, however, declined from .757.757.757.757 to .527.527.527.527 during the study period when the trade system shifted from less to more multipolar.
Figure V: The sum of $\pi_{i}$ for all other countries (2000-2019).
Figure VI: Correlation between $\pi_{i}$ and $g^{2}_{i}$ after controlling $g_{i}$ (2000-2019).
Table I: Country $j$’s direct contribution to $\pi_{i}$ for 2017 and 2019 (in percentage)
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA $\pi_{i}$
CAN $\frac{76.647}{77.390}$ $\frac{1.2635}{1.2732}$ $\frac{0.5390}{0.5179}$ $\frac{0.3222}{0.2781}$ $\frac{0.5922}{0.5728}$ $\frac{0.2877}{0.2607}$ $\frac{0.4329}{0.4768}$ $\frac{0.0410}{0.0355}$ $\frac{14.492}{13.916}$$\frac{.025965}{.026125}$
CHN $\frac{0.3952}{0.3439}$ $\frac{85.584}{84.126}$ $\frac{0.8233}{0.8565}$ $\frac{0.3967}{0.3931}$ $\frac{0.3423}{0.3479}$ $\frac{0.1878}{0.2052}$ $\frac{1.3131}{1.4868}$ $\frac{0.3110}{0.3095}$ $\frac{1.5849}{2.1950}$$\frac{.093905}{.086596}$
DEU $\frac{0.2088}{0.1682}$ $\frac{1.2344}{1.2456}$ $\frac{67.501}{67.410}$ $\frac{2.2469}{2.3993}$ $\frac{1.8879}{2.0095}$ $\frac{1.4928}{1.5634}$ $\frac{0.4265}{0.3507}$ $\frac{0.3737}{0.5379}$ $\frac{3.2504}{3.0282}$$\frac{.040521}{.040105}$
FRA $\frac{0.2345}{0.2171}$ $\frac{0.7032}{0.6617}$ $\frac{3.5649}{3.5687}$ $\frac{74.940}{75.190}$ $\frac{1.9053}{1.8352}$ $\frac{1.6149}{1.5794}$ $\frac{0.2216}{0.2111}$ $\frac{0.1656}{0.2247}$ $\frac{2.4124}{2.3005}$$\frac{.030196}{.030772}$
GBR $\frac{0.6234}{0.6356}$ $\frac{0.9550}{0.9428}$ $\frac{2.2098}{2.4071}$ $\frac{1.5773}{1.6503}$ $\frac{75.873}{75.862}$ $\frac{0.6921}{0.7259}$ $\frac{0.3716}{0.4077}$ $\frac{0.1986}{0.1736}$ $\frac{4.0273}{3.5605}$$\frac{.037385}{.036390}$
ITA $\frac{0.1418}{0.1419}$ $\frac{1.0159}{0.9305}$ $\frac{3.2895}{3.1651}$ $\frac{2.3107}{2.3567}$ $\frac{1.3228}{1.2695}$ $\frac{74.945}{75.662}$ $\frac{0.1958}{0.2123}$ $\frac{0.4274}{0.4525}$ $\frac{1.7055}{1.5402}$$\frac{.019532}{.020507}$
JPN $\frac{0.3297}{0.3118}$ $\frac{2.0275}{1.9912}$ $\frac{0.5242}{0.4869}$ $\frac{0.3184}{0.3041}$ $\frac{0.4209}{0.3614}$ $\frac{0.2218}{0.2139}$ $\frac{87.560}{87.198}$ $\frac{0.1581}{0.1924}$ $\frac{2.2315}{1.9753}$$\frac{.043971}{.045859}$
RUS $\frac{0.1182}{0.1099}$ $\frac{2.6727}{2.2351}$ $\frac{1.7330}{1.7824}$ $\frac{0.9775}{1.0709}$ $\frac{0.7020}{0.7266}$ $\frac{0.8945}{0.8129}$ $\frac{0.6044}{0.5044}$ $\frac{80.225}{79.511}$ $\frac{1.5211}{1.6313}$$\frac{.010657}{.012509}$
USA $\frac{1.3878}{1.4971}$ $\frac{0.8107}{1.0322}$ $\frac{0.4527}{0.4749}$ $\frac{0.2580}{0.2740}$ $\frac{0.5935}{0.5643}$ $\frac{0.1733}{0.1802}$ $\frac{0.3514}{0.3897}$ $\frac{0.0368}{0.0464}$ $\frac{92.142}{91.369}$$\frac{.293880}{.272142}$
* The numerators are for 2019 and the denominators for 2017.
A straightforward implication of Eq. (4) is country $j$’s direct contribution to country $i$’s competitiveness in percentage, i.e., $\pi_{j}P_{ji}/\pi_{i}\times 100\%$. Table I lists these contributions for 2017 and 2019. According to the diagonal numbers, the USA was the most self-sufficient in feeding its competitiveness, with a slight increase from 2017 to 2019. The ninth column also shows that the USA was the most significant contributor to Canada, China, Germany, Japan, and the UK in 2019. For the same year, China was the most significant contributor to Russia’s competitiveness and the second largest to those of Canada, Japan, and the USA. France, Germany, and Italy were highly reliant on each other’s contributions, and their average dependence on Russia was 8.7 times higher than the USA’s dependence on Russia. In value, the USA contributed .14492.025965=.003763.14492.025965.003763.14492*.025965=.003763.14492 ∗ .025965 = .003763 to Canada while Canada contributed .013878.293880=.004078.013878.293880.004078.013878*.293880=.004078.013878 ∗ .293880 = .004078 to the USA in 2019. Comparing the years before and after the beginning of the China-USA trade war, the USA’s contributions to China dropped by $21.7\%$, calculated as 1(1.5849.093905)/(2.1950.086596)11.5849.0939052.1950.0865961-(1.5849*.093905)/(2.1950*.086596)1 - ( 1.5849 ∗ .093905 ) / ( 2.1950 ∗ .086596 ), while China’s contribution to the USA reduced by $15.2\%$, calculated as 1(.8107.29388)/(1.0322.272142)1.8107.293881.0322.2721421-(.8107*.29388)/(1.0322*.272142)1 - ( .8107 ∗ .29388 ) / ( 1.0322 ∗ .272142 ).
Figure VII: Bargaining power $\beta_{i}$ (2000-2019).
Figure VII illustrates the national bargaining power from 2000 to 2019, displaying patterns similar to those in Figure IV, albeit smoother. China’s bargaining power steadily increased due to rapid production growth, even during its $\pi_{i}$’s troughs in 2005 and 2015. Notably, China’s bargaining power consistently surpassed its competitiveness, whereas the USA’s bargaining power was consistently lower than its competitiveness, indicating $\beta$’s mitigation of the Matthew effect in $\pi$. Russia’s bargaining power sharply declined following the Russo-Georgian War in 2008 and the annexation of Crimea in 2014.
Figure VIII illustrates the comparative bargaining power of the USA against other countries. The bargaining power of four Western European countries experienced a slight increase following the establishment of the Eurozone in 1999, but saw a modest decline post-2008, culminating in an average loss of 15.7% of their bargaining power to the USA. Conversely, China and Russia significantly enhanced their relative power multiple times within the first 15 years, maintaining a steady level thereafter. Japan, however, suffered a substantial reduction, losing approximately 50% of its relative bargaining power. Canada’s relative bargaining strength surged during crises but swiftly reverted to its original level, attributable to its strong alignment with the USA.
Figure VIII: Bargaining power versus the USA (2000-2019).
Table II: $1,000\times\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{-25.78}{-5.78}$ $\frac{-5.49}{-0.37}$ $\frac{-1.47}{0.29}$ $\frac{-0.72}{2.07}$ $\frac{-2.35}{-0.16}$ $\frac{-4.87}{-5.35}$ $\frac{-0.62}{-0.53}$ $\frac{15.75}{63.78}$
CHN $\frac{6.58}{6.90}$ $\frac{7.39}{16.62}$ $\frac{5.36}{7.76}$ $\frac{6.05}{8.46}$ $\frac{2.14}{5.25}$ $\frac{11.56}{31.43}$ $\frac{0.89}{-2.48}$ $\frac{45.19}{65.48}$
DEU $\frac{1.18}{0.14}$ $\frac{-7.73}{-6.05}$ $\frac{6.49}{3.43}$ $\frac{10.07}{5.30}$ $\frac{1.29}{0.18}$ $\frac{0.10}{-5.13}$ $\frac{-1.86}{-4.59}$ $\frac{23.54}{15.06}$
FRA $\frac{0.98}{-0.19}$ $\frac{-8.51}{-5.16}$ $\frac{-13.75}{-6.39}$ $\frac{5.54}{2.67}$ $\frac{-4.20}{-2.57}$ $\frac{-0.90}{-3.67}$ $\frac{-2.08}{-3.10}$ $\frac{12.90}{9.72}$
GBR $\frac{0.73}{-0.96}$ $\frac{-16.47}{-7.66}$ $\frac{-18.11}{-7.01}$ $\frac{-5.31}{-2.28}$ $\frac{-4.64}{-2.58}$ $\frac{-3.39}{-7.65}$ $\frac{-2.38}{-1.99}$ $\frac{8.25}{10.81}$
ITA $\frac{1.01}{0.14}$ $\frac{-8.29}{-5.79}$ $\frac{-3.66}{-0.48}$ $\frac{6.12}{3.66}$ $\frac{7.17}{4.41}$ $\frac{-0.21}{-2.97}$ $\frac{-2.65}{-6.84}$ $\frac{12.51}{11.97}$
JPN $\frac{4.31}{3.52}$ $\frac{-26.28}{-14.31}$ $\frac{-0.28}{4.76}$ $\frac{2.01}{3.72}$ $\frac{4.33}{6.66}$ $\frac{0.35}{2.27}$ $\frac{-1.77}{-2.08}$ $\frac{30.88}{69.74}$
RUS $\frac{1.52}{4.49}$ $\frac{-5.95}{3.48}$ $\frac{8.91}{56.17}$ $\frac{9.85}{24.14}$ $\frac{9.05}{26.41}$ $\frac{5.10}{25.58}$ $\frac{2.78}{6.37}$ $\frac{24.25}{56.37}$
USA $\frac{-1.78}{-5.56}$ $\frac{-48.89}{-17.13}$ $\frac{-12.58}{-6.53}$ $\frac{-4.11}{-2.91}$ $\frac{-3.01}{-2.77}$ $\frac{-4.26}{-2.95}$ $\frac{-10.14}{-23.53}$ $\frac{-2.18}{-1.84}$
* The numerators are for 2017 and the denominators for 2000.
Table II presents the derivatives of $\log\pi_{i}$ with respect to $\log P_{ji}$ for the years 2000 and 2017, calculated from Eqs. (9) and (12). For instance, a 1% reduction in imports from China by the USA in 2017 would have resulted in an increase of .04889% in the USA’s competitiveness. Although the USA may have had reasons to initiate trade wars against all eight other countries, its primary targets were China in 2017 and Japan in 2000, as indicated by the numerators and denominators in the last row. As evidenced by the last column, China and Japan would have been the two largest beneficiaries had they further collaborated with the USA in 2017 and 2000, respectively. However, a comparison of the numerators and denominators in the column reveals a considerable reduction in benefits from 2000 to 2017. Additionally, initiating trade friction from any of these eight countries against the USA would result in a loss for the initiating country. A trade war against Russia would be nearly fruitless due to the small numerators and denominators in the eighth column. Furthermore, the numbers symmetric with the diagonal exhibit opposite signs, corroborating Corollary 3.
Table III: $1,000\times\frac{\mathrm{d}\log\pi_{j}}{\mathrm{d}\log P_{ii}}$ for years 2000 and 2017
${}_{i}\mkern-6.0mu\setminus\mkern-6.0mu{}^{j}$
CAN CHN DEU FRA GBR ITA JPN RUS USA
CAN $\frac{52.87}{-188.64}$ $\frac{-16.74}{-9.79}$ $\frac{5.29}{17.08}$ $\frac{4.44}{17.11}$ $\frac{-6.47}{13.46}$ $\frac{5.43}{15.50}$ $\frac{-9.91}{3.11}$ $\frac{3.34}{4.84}$ $\frac{-10.59}{-10.07}$
CHN $\frac{195.55}{35.76}$ $\frac{-2324.27}{-1564.33}$ $\frac{106.47}{31.54}$ $\frac{180.96}{36.26}$ $\frac{232.75}{40.17}$ $\frac{126.87}{31.74}$ $\frac{-81.64}{-33.03}$ $\frac{-135.91}{-2.35}$ $\frac{186.27}{38.81}$
DEU $\frac{38.68}{39.59}$ $\frac{-10.75}{-52.20}$ $\frac{-253.88}{-139.10}$ $\frac{-38.50}{-44.43}$ $\frac{0.50}{-0.04}$ $\frac{-45.19}{-47.50}$ $\frac{14.09}{7.10}$ $\frac{-37.45}{-205.30}$ $\frac{41.38}{44.19}$
FRA $\frac{21.86}{20.97}$ $\frac{-9.83}{-22.47}$ $\frac{-69.45}{-62.18}$ $\frac{-41.92}{-12.66}$ $\frac{-17.07}{-22.86}$ $\frac{-67.71}{-56.16}$ $\frac{1.87}{1.19}$ $\frac{-43.68}{-97.76}$ $\frac{22.93}{25.18}$
GBR $\frac{15.92}{-5.81}$ $\frac{-12.59}{-23.38}$ $\frac{-67.31}{-54.42}$ $\frac{-49.89}{-44.88}$ $\frac{164.52}{145.52}$ $\frac{-42.47}{-29.07}$ $\frac{-7.26}{-13.58}$ $\frac{-33.94}{-117.05}$ $\frac{17.78}{26.54}$
ITA $\frac{18.26}{19.31}$ $\frac{1.06}{-15.52}$ $\frac{-48.34}{-45.02}$ $\frac{-36.50}{-36.91}$ $\frac{-1.16}{-0.96}$ $\frac{-257.39}{-190.00}$ $\frac{4.94}{0.75}$ $\frac{-32.92}{-135.50}$ $\frac{20.83}{23.82}$
JPN $\frac{66.38}{212.93}$ $\frac{-191.11}{-612.51}$ $\frac{35.78}{163.77}$ $\frac{56.09}{212.97}$ $\frac{61.37}{201.14}$ $\frac{46.13}{200.28}$ $\frac{-726.94}{-1518.40}$ $\frac{-19.37}{117.48}$ $\frac{69.31}{236.68}$
RUS $\frac{12.12}{7.85}$ $\frac{-20.29}{-11.83}$ $\frac{-16.42}{4.56}$ $\frac{4.27}{7.57}$ $\frac{8.86}{8.30}$ $\frac{-3.98}{3.36}$ $\frac{-3.65}{3.75}$ $\frac{-832.77}{-2139.12}$ $\frac{12.34}{8.00}$
USA $\frac{-142.03}{-638.28}$ $\frac{-567.12}{-794.20}$ $\frac{-449.08}{-239.99}$ $\frac{-308.90}{-168.36}$ $\frac{-283.70}{-260.03}$ $\frac{-279.33}{-167.92}$ $\frac{-411.84}{-742.34}$ $\frac{-410.17}{-643.02}$ $\frac{834.90}{640.50}$
* The numerators are for 2017 and the denominators are for 2000.
Table III lists the derivatives of $\log\pi$ with respect to $\log P_{ii}$ for 2017 and 2000, calculated from Eqs. (17), (18), and (12). From the diagonal, globalization significantly benefited China, Germany, Italy, Japan, and Russia. It was advantageous for Canada in 2000 but detrimental in 2017. The effects on France were insignificant due to small numerator and denominator. The USA and the UK would be anti-globalization advocates, explaining recent events such as the Trump administration’s withdrawal from international organizations and the UK’s Brexit. Though business interest groups overwhelmingly profited from the gigantic international markets and low-priced raw materials, growing low-income Americans from manufacturing offshoring demanded more inexpensive goods from China, uplifting trade deficits and national debts but benefiting from low inflation and affordable merchandises. Also, China was the largest beneficiary of globalization in 2017 with a 2.324% hike in $\pi_{i}$ for a 1% decrease in $P_{ii}$, so its Belt and Road Initiative (BRI) would have a positive impulse on its $\pi_{i}$. Russia was the largest beneficiary of globalization in 2000.
The side effects on third parties are positioned off the diagonal. For instance, the USA’s further globalization benefited all eight other countries, thereby making its protectionist policies likely to face significant resistance from its partners. China emerged as the largest beneficiary of the USA’s globalization in both years. Conversely, Canada experienced the least impact from the USA’s protectionism in 2017; its further globalization would have marginally benefited the USA, as indicated by the first row. The last column suggests that further globalization from seven other countries would adversely affect the USA; Japan and China topped the list in 2000 and 2017, respectively. Their adverse impact decreased by 70.2% and increased by 380%, respectively, over the 17 years. Additionally, the second row elucidates the USA and its Western allies’ objection to China’s BRI-like projects, which intensified by $4.74$ times from 2000 to 2017, according to the ratios of their numerators and denominators.
We assess the externalities on the third parties due to the USA’s trade war against China in 2018 and 2019. Table IV is derived from Eqs. (9), (11), and (12). When the USA reduced its imports from China by 1% in 2018 and 2019, China’s $\pi_{j}$ suffered a significant loss of .1058% and .0937%, respectively, while the USA’s $\pi_{i}$ earned a non-significant gain of .0540% and .0460%, respectively. Canada was a free rider, and the other six countries were slightly negatively affected in 2018 and 2019. Consequently, these countries would likely refrain from taking formal action to prevent the trade war. Furthermore, Canada or Russia would be the best ally with the USA or China, respectively, in this conflict. The second-best allies would be the UK for the USA and Japan for China, respectively. From 2018 to 2019, the effects were slightly mitigated for all these countries. Likely recognizing the mutually destructive consequences of the trade war, China and the USA agreed on a new trade deal in early 2020 (cf, US Trade Representative, 2020), concluding the first phase of a long-term conflict. However, the cumulative effects would be significant as the conflict is expected to continue for years.
Table IV: $1000\times\frac{\mathrm{d}\log\pi_{k}}{\mathrm{d}\log P_{ji}}$ for the China-USA Trade War’s side effects on third parties
Year CAN DEU FRA GBR ITA JPN RUS CHN USA
2018 -26.114 7.103 3.013 .444 6.661 13.639 15.097 105.808 -53.960
2019 -21.815 3.223 2.133 .039 5.324 12.392 14.894 93.744 -45.954
* Country $i$ is for the USA, $j$ for China, and $k$ for others.
Finally, the cooperative solution $\lambda_{ji}^{*}$ could mitigate the conflict and prevent further escalation. At the inception of this conflict in 2018,
$\frac{\pi_{j}}{\pi_{i}}=\frac{.092618}{.284327}<\frac{g_{j}}{g_{i}}=\frac{17494.79}{23511.17}.$
For a positive $\mathrm{d}P_{ji}$, any $\lambda_{ji}$ between $\pi_{j}/\pi_{i}$ and $g_{j}/g_{i}$ would not only make $\frac{\mathrm{d}\log\pi_{i}}{\mathrm{d}\log P_{ji}}<0$ but also create further trade deficits for the USA (see Table V). Thus, an economic conflict with a negative $\mathrm{d}P_{ji}$ was possible and, in this lose-lose conflict, the Nash bargaining solution was:
[…]
for 2018. For each dollar’s decrease in exports from China to the USA, $\mathrm{d}P_{ji}=-\$1/g_{j}$. According to the bargaining solution, China would have changed its imports from the USA by $\mathrm{d}P_{ij}=\lambda_{ji}^{*}\mathrm{d}P_{ji}$ of the USA’s production, which is .$66164.49233*(-1/g_{j})*g_{i}=-\$.$66164. Therefore, if this resolution had been applied, the USA would have earned \$.33836=(-.66164)-(-1) surplus for each decreased dollar of imports from China. Additionally, by Theorem  2, the USA’s $\pi_{i}$ would have increased since $\lambda^{*}_{ji}>\pi_{j}/\pi_{i}$ and $\mathrm{d}P_{ji}<0$. To maintain its competitive advantage $\pi_{i}$ and reduce national debt by earning a trade surplus in the bilateral friction, by the impossibility trilemma in Theorem 4 , the USA had to sacrifice its exports and shrink its imports — in either case, $\mathrm{d}P_{ij}<0$.
7 Discussions
Economic globalization and trade wars have emerged as significant public interests due to global competition among nations, large-scale restructuring of labor markets, and the international distribution of added value. This study posits that import and export data already reflect these issues, and that competitive advantages, rather than comparative advantages, are the primary drivers of trade wars and de-globalization. Consequently, we aim to extract a nation’s competitive strength and bargaining power from the data, identify optimal strategies for addressing trade conflicts and cooperation, and determine the extent to which a nation should pursue globalization.
The methodologies employed, grounded in network and game theories, focus on contemporaneously interactive relationships among economies and are devoid of parameters and pre-set utility functions. These network game approaches can be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
Several critical factors warrant further investigation. First, counterbalance disequilibrium may arise when trades are notably imbalanced. This imbalance can manifest in various ways. For instance, when country $i$ uses borrowed money to consume country $j$’s exports, country $i$’s consumption still exerts influence over country $j$’s production. However, this influence originates not only from country $i$, but also from the money lender, who possesses some control over $i$’s consumption. If the lender happens to be country $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ is underestimated, potentially leading to a debt trap for country $i$.
Secondly, unlike an economic superpower, an emerging economy may prioritize the growth of exports over competitive advantages. Indeed, its economic growth must significantly surpass that of advanced economies to bridge the competitiveness gap, in accordance with $\pi$’s Matthew effect. Thus, maximizing competitive advantages may be a low priority for an emerging economy. For example, during China’s primitive capital accumulation in the 1980s and 1990s, earning additional US dollars was essential to finance its industrial modernization. The time dimension, absent from the current research, is necessary to formulate growth strategies.
Thirdly, the facets of economic globalization encompassing capital, information, and technology have yet to be fully integrated. These elements interact with goods and services, but their data are not as easily accessible. Additionally, several data issues may undermine the empirical study’s results. Notably, there are numerous missing values in the service sectors, and discrepancies often arise between country $i$’s reported imports from country $j$ and country $j$’s reported exports to country $i$.
Furthermore, the integration of regional economies warrants further examination. Theorem 2 could be employed to identify suitable candidates for integration. However, the underlying assumption may need modification. For instance, if country $i$ is a member of multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries as used in Theorem 6.
Moreover, additional variables could be incorporated into the demand-sided counterbalance Eq. (3), particularly those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. For example, incorporating labor loss into $g_{i}$. Any factors directly influencing country $i$’s production could be included in the $i$th row of $P$. Although consumption may still predominantly drive the fractions, not all production is driven by consumption.
The framework can be extended in several directions. Firstly, amid a bilateral trade war and anti-globalization sentiments, a country may target specific groups of partners without directly impacting others. For instance, as illustrated in Table II, the USA could simultaneously increase tariffs on imports from China, Germany, and Japan.
Secondly, while the dynamical system described by Eqs. (9)–(11) or Eqs. (17)–(18) assumes a fixed $P$, introducing a temporal dimension to $P$ could provide valuable insights. Incorporating a time horizon would facilitate the examination of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and the improvement of trade deficits. Additionally, the temporal dimension could enhance the model’s forecast capabilities. Intertemporal delays often occur when production takes place in one year but consumption in another. If production spans two years, the previous year’s imports of intermediate goods could be included in the current year’s completed production.
Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios. An industrial or geopolitical analysis could provide valuable insights into this choice. For instance, as Japan and South Korea competed in automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was appropriate for $\lambda_{ji}$. In the same year, no industry served as a battleground for China and the USA, making $g_{j}/g_{i}$ a natural choice for $\lambda_{ji}$. Moreover, the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the importance of the distance between trade partners and the type of merchandise as crucial determinants of trade flows. Consequently, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
Econometric analyses could also provide additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is greater or less than 00, we may infer that $\pi_{i}$ is over- or under-valued, respectively. Including additional regressors, such as debt to GDP ratio, in the regression helps to better understand $\pi$ through explanatory data.
Finally, anticipating the side effects of country $i$’s trade war against country $j$, country $j$ might seek cooperation with third parties that would substantially suffer from the conflict. However, adding more countries to both sides could result in a global trade war, akin to the Cold War between 1947 and 1991.
Appendix
A1. Proof of Theorem 1
We multiply $\vec{1}_{n}$ from the right on both sides of the equation $\rho P=c\rho$ to get $\rho P\vec{1}_{n}=c\rho\vec{1}_{n}.$ Using $P\vec{1}_{n}=\vec{1}_{n}$ and $\rho\vec{1}_{n}=1$, we obtain $\rho\vec{1}_{n}=c.$ Thus, $c=1$. Also, by $\rho P=c\rho=\rho$ and the uniqueness of $\pi$, we have $\rho=\pi$.
A2. Proof of Corollary 1
Assume that $g_{j}>g_{k}$. For example, country $i$ considers certain imports from either country $j$ or $k$, all else being equal. If $i$ selects $j$, by Eq. (4), then the competitiveness country $i$ derives from $j$ in the imports is approximately
$\pi_{j}\Delta P_{ji}=\pi_{j}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j}{g_{j}}=\frac{\pi_{j}}{g_{j}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ j)$
where $\Delta P_{ji}$ is the change on $P_{ji}$ due to the imports. The approximation ignores the ripple effect from the shock. If $i$ selects $k$, then the competitiveness country $i$ derives from $k$ in the imports is approximately
$\pi_{k}\Delta P_{ki}=\pi_{k}\times\frac{\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k}{g_{k}}=\frac{\pi_{k}}{g_{k}}\times(\mathrm{the}\ \mathrm{imports}\ \mathrm{from}\ k).$
Because of the Matthew effect, ${\pi_{j}}/{g_{j}}>{\pi_{k}}/{g_{k}}$, and so, country $i$ would likely choose $j$, all else being equal.
A3. Proof of Theorem 2
We apply matrix calculus with the restrictions of Eq. (3) and $\pi\vec{1}_{n}=1$ to prove the theorem. When making a small shock or perturbation $\Delta P$ to $P$ in Eq. (3), the new authority distribution $\pi+\Delta\pi$ satisfies the counterbalance equation:
$\pi+\Delta\pi=(\pi+\Delta\pi)[P+\Delta P]$ (A.1)
subject to $\Delta P\vec{1}_{n}=\vec{0}_{n}$ and $\Delta\pi\vec{1}_{n}=0$. After subtracting $\pi=\pi P$ from Eq. (A.1), we get $\Delta\pi[I_{n}-P]=\pi\Delta P+\Delta\pi\Delta P$ and its first-order approximation is:
$\Delta\pi[I_{n}-P]\approx\pi\Delta P.$ (A.2)
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$. Accordingly, $P$ has three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$. Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\left[I_{n}-P\right]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\left(\vec{0}^{\top}_{i-1},-\lambda_{ji}\pi_{i}+\pi_{j},\vec{0}^{\top}_{j-i-1},\lambda_{ji}\pi_{i}-\pi_{j},\vec{0}^{\top}_{n-j}\right)$ (A.3)
where the right-hand vector has all zeros except for the $i$th and the $j$th elements.
Also, we partition the transpose of $P$ as
$P^{\top}=\left[\begin{array}[]{ccccc}H_{1}&\eta_{1}&H_{2}&\eta_{2}&H_{3}\\ \mu_{1}&\boxed{P_{ii}}&\mu_{2}&\boxed{P_{ji}}&\mu_{3}\\ H_{4}&\eta_{3}&H_{5}&\eta_{4}&H_{6}\\ \mu_{4}&\boxed{P_{ij}}&\mu_{5}&\boxed{P_{jj}}&\mu_{6}\\ H_{7}&\eta_{5}&H_{8}&\eta_{6}&H_{9}\end{array}\right]$ (A.4)
where $\eta_{1},\cdots,\eta_{6}$ are column vectors, $\mu_{1},\cdots,\mu_{6}$ are row vectors, and $H_{1},\cdots,H_{9}$ are sub-matrices of $P^{\top}$. We write the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ji}}\vec{1}_{n}=0$ and the transpose of Eq. (A.3) as
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&\vec{0}_{i-1}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&-\lambda_{ji}\pi_{i}+\pi_{j}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&\vec{0}_{j-i-1}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\lambda_{ji}\pi_{i}-\pi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&\vec{0}_{n-j}\end{array}\right].$
Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the $(i+1)$st row, making the $(i+1)$st row of Eq. (A.5) a zero vector. After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$ (A.6)
To make the matrix in Eq. (A.6) lower-triangular, we multiply the following non-singular matrix
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.7)
to the left side of Eq. (A.6) to get:
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.8):
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}=-\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.8):
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
and thus:
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}\right]\\ &=&(\lambda_{ji}\pi_{i}-\pi_{j})\left(I_{n-1}-Z_{i}\right)^{-1}\left[\gamma_{ji}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\gamma_{ji}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
Similarly, after we add all rows (except for the first one) in Eq. (A.5) to the $(j+1)$st row, the $(j+1)$st row becomes a zero vector. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$ (A.9)
We multiply the following non-singular matrix:
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
to the left side of Eq. (A.9) to get:
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}&\vec{0}_{n-1}^{\top}&(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}\\ -\alpha_{j}&I_{n-1}-Z_{j}&(-\lambda_{ji}\pi_{i}+\pi_{j})\gamma_{ij}\\ \end{array}\right].$
Therefore, by the first row of Eq. (A.10),
$\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}=\frac{(\lambda_{ji}\pi_{i}-\pi_{j})\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{j})^{-1}\gamma_{ij}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\alpha_{j}}.$
A4. Proof of Corollary 3
First, both $(I_{n-1}-Z_{i})^{-1}=I_{n-1}+Z_{i}+Z_{i}^{2}+Z_{i}^{3}+\cdots$ and $(I_{n-1}-Z_{j})^{-1}=I_{n-1}+Z_{j}+Z_{j}^{2}+Z_{j}^{3}+\cdots$ are non-negative matrices. Secondly, $\vec{1}_{n-1}$, $\gamma_{ji}$, $\gamma_{ij}$, $\alpha_{i}$, and $\alpha_{j}$ are all non-negative. Therefore, by Eqs. (9) and (10) in Theorem 2, $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0$. Also,
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ij}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ij}}=\frac{\mathrm{d}\pi_{i}}{\lambda_{ji}\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\lambda_{ji}\mathrm{d}P_{ji}}=\frac{1}{\lambda_{ji}^{2}}\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}}\leq 0.$
A5. Proof of Theorem 3
In added value, the changed exports from country $j$ to $i$ are $g_{j}\Delta P_{ji}$, while those from country $i$ to $j$ are $g_{i}\Delta P_{ij}=g_{i}\lambda_{ji}\Delta P_{ji}$. If $\lambda_{ji}=g_{j}/g_{i}$, then $g_{j}\Delta P_{ji}=g_{i}\Delta P_{ij}$, resulting in zero net trade surplus and deficit between these countries. This condition is for instant balance change.
Over a specific time period, there is no trade surplus or deficit if and only if $g_{j}/g_{i}=P_{ij}/P_{ji}$, because $g_{i}P_{ij}$ represents exports from $i$ to $j$ and $g_{j}P_{ji}$ represents exports from $j$ to $i$.
A6. Proof of Theorem 4
When $\pi_{j}/\pi_{i}<g_{j}/g_{i}$, there are three possible intervals and two values for $\lambda_{ji}$: $0<\lambda_{ji}<\pi_{j}/\pi_{i}$, $\lambda_{ji}=\pi_{j}/\pi_{i}$, $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$, $\lambda_{ji}=g_{j}/g_{i}$, or $g_{j}/g_{i}<\lambda_{ji}<\infty$. The small change $\Delta P_{ji}$ could be positive or negative, and $\Delta P_{ij}=\lambda_{ji}\Delta P_{ji}$ has the same sign. Using Theorems 2 and 3, Table V lists all possible outcomes for these ten scenarios of $(\lambda_{ji},\Delta P_{ji})$.
Table V: Scenarios of $(\lambda_{ji},\Delta P_{ji})$ and their impacts on country $i$’s $\pi_{i}$ and net trade with $j$
$0<\lambda_{ji}<\frac{\pi_{j}}{\pi_{i}}$ $\lambda_{ji}=\frac{\pi_{j}}{\pi_{i}}$ $\frac{\pi_{j}}{\pi_{i}}<\lambda_{ji}<\frac{g_{j}}{g_{i}}$ $\lambda_{ji}=\frac{g_{j}}{g_{i}}$ $\frac{g_{j}}{g_{i}}<\lambda_{ji}<\infty$
$\Delta P_{ji}>0$ $\pi_{i}\uparrow$ $\Delta\pi_{i}=0$ $\pi_{i}\downarrow$ $\pi_{i}\downarrow\downarrow$ $\pi_{i}\downarrow\downarrow\downarrow$
---- - - --- - -- zero net $+$
$\Delta P_{ji}<0$ $\pi_{i}\downarrow$ $\Delta\pi_{i}=0$ $\pi_{i}\uparrow$ $\pi_{i}\uparrow\uparrow$ $\pi_{i}\uparrow\uparrow\uparrow$
+++++++ + + +++++ + $+$ zero net --
$+$ and -- for net trade surplus and deficit, respectively.
Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, $+$, or --.
$\Delta\pi_{j}$ and $\Delta\pi_{i}$ have opposite signs and country $i$ and $j$’s net trade balances have opposite signs.
From the first row, country $i$ could not increase $\pi_{i}$ and achieve a trade surplus with country $j$ at the same time. When $\pi_{j}/\pi_{i}<\lambda_{ji}<g_{j}/g_{i}$ in the same row, however, country $j$ has a trade surplus, an increasing $\pi_{j}$, and $\Delta P_{ji}>0$. In these ten scenarios, none strictly dominates the others, regarding competitiveness, net trade balance, comparative advantages, and their magnitudes of changes. On the other hand, countries $i$ and $j$ may not completely contradict in all these interests.
A7. Proof of Theorem 5
In the first scenario, country $j$ prefers $q$ to $p$, whereas country $i$ prefers $1/p$ to $1/q$. Therefore, $u_{j}(x)=x$ is a utility function for $j$ because $u_{j}(q)>u_{j}(p)$. Similarly, $u_{i}(x)=x$ is a utility function for $i$. These are von Neumann-Morgenstern utilities, uniquely determined up to a positive affine transformation. Assuming that all countries have equal prior bargaining power, the Nash bargaining solution solves the following maximization problem:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-p\right)\left(\lambda_{ij}-\frac{1}{q}\right).$ (A.11)
In Eq. (A.11), $\lambda_{ji}-p$ and $\lambda_{ij}-\frac{1}{q}$ are the excessive payoffs or welfare for countries $j$ and $i$ in terms of their respective utility functions. The bargaining solution maximizes the product of the excessive utilities.
In the second scenario of opposite preferences, country $j$ prefers $p$ to $q$ while country $i$ prefers $1/q$ to $1/p$. Their utility functions could be $u_{j}(x)=u_{i}(x)=-x$, and their status quo points are $u_{j}(q)=-q$ and $u_{i}(1/p)=-1/p$, respectively. The maximization problem becomes:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(-\lambda_{ji}+q\right)\left(-\lambda_{ij}+\frac{1}{p}\right).$ (A.12)
In this utility function, $\lambda_{ji}$ seems to have no meaning of reprisal. However, we can rephrase Eq. (A.12) by:
$\max\limits_{\lambda_{ji}\lambda_{ij}=1;\ \lambda_{ji}>0}\left(\lambda_{ji}-q\right)\left(\lambda_{ij}-\frac{1}{p}\right)$ (A.13)
in which $\lambda_{ji}-q$ and $\lambda_{ij}-1/p$ are the loss functions for $j$ and $i$, respectively. To coerce an agreement formation, each country plays a threat strategy to maximize its counterpart’s loss. The final outcome is Eq. (A.13). Moving from $p$ to $q$, a larger $\lambda_{ji}$ means less loss to $j$; thus, $\lambda_{ji}$ still preserves the meaning of retaliation for $j$ in Eq. (A.13) and, so, in Eq. (A.12).
Let $x=\lambda_{ji}$. Then $\lambda_{ij}=1/x$. By Eq. (A.11), the Nash bargaining solution is:
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-p\right)\left(\frac{1}{x}-\frac{1}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{q}+\frac{p}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Also, by Eq. (A.12) or (A.13):
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-q\right)\left(\frac{1}{x}-\frac{1}{p}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{x}{p}+\frac{q}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
A8. Proof of Corollary 4
When we extend the effective bargaining set of $\lambda_{ji}$ from $[p,q]$ to $[cp,q/c]$, the corresponding range for $\lambda_{ij}$ then changes from $[1/q,1/p]$ to $[c/q,1/(cp)]$. Letting $x=\lambda_{ji}$, the bargaining solution to Eq. (A.11) becomes:
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(x-cp\right)\left(\frac{1}{x}-\frac{c}{q}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{cx}{q}+\frac{cp}{x}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
In the above, we need $c>0$ and $cp\leq q/c$, i.e., $0<c\leq\sqrt{q/p}$.
After we replace $p$ and $q$ with $cp$ and $q/c$, respectively, Eq. (A.12) or (A.13) becomes:
$\mathop{\mathrm{argmax}}\limits_{x>0}\left(\frac{q}{c}-x\right)\left(\frac{1}{cp}-\frac{1}{x}\right)=\mathop{\mathrm{argmin}}\limits_{x>0}\left(\frac{q}{cx}+\frac{x}{cp}\right)=\sqrt{pq}=\sqrt{\frac{\pi_{j}g_{j}}{\pi_{i}g_{i}}}.$
Finally, it is easy to check that $\sqrt{pq}\in[cp,q/c]$ for any $0<c\leq\sqrt{q/p}$.
A9. Proof of Corollary 5
This is because:
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{\pi_{j}}{\pi_{i}}\right)\left(\lambda_{ij}-\frac{\pi_{i}}{\pi_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{\pi_{j}}{\pi_{i}}$ (A.14)
and
$\underset{\lambda_{ji}>0}{\mathrm{argmax}}\left\{\left(\lambda_{ji}-\frac{g_{j}}{g_{i}}\right)\left(\lambda_{ij}-\frac{g_{i}}{g_{j}}\right)\bigg{|}\lambda_{ij}\lambda_{ji}=1\right\}=\frac{g_{j}}{g_{i}}.$ (A.15)
In Eqs. (A.14) and (A.15), both countries battle for better competitiveness and net trade balance, respectively. At the solution points, the objective functions are zero and no country gets extra welfare,
A10. Proof of Corollary 6
Country $j$’s direct competition with $i$ for a higher $\pi_{j}$ results in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, regardless of trade surplus or deficit.
Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$.
A11. Proof of Theorem 6
We write $M=[\xi_{1},\xi_{2},\cdots,\xi_{n}]$ where $\xi_{j}$ is the $j$th column of $M$. Since $\sigma_{ji}$ has a zero sum for all $j$, $\xi_{1}+\xi_{2}+\cdots+\xi_{n}=\vec{0}_{n}$. We divide Eq. (A.2) by $\Delta P_{ii}$ and let $\Delta P_{ii}\to 0$ to get:
$\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}[I_{n}-P]=\pi\frac{\mathrm{d}P}{\mathrm{d}P_{ii}}=\pi M=\left(\pi\xi_{1},\cdots,\pi\xi_{n}\right).$ (A.16)
With the partition Eq. (A.4) of $P^{\top}$, the augmented matrix for the identity $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}\vec{1}_{n}=0$ and the transpose of Eq. (A.16) is:
$\left[\begin{array}[]{ccccc|c}\vec{1}_{i-1}^{\top}&1&\vec{1}_{j-i-1}^{\top}&1&\vec{1}_{n-j}^{\top}&0\\ I_{i-1}-H_{1}&-\eta_{1}&-H_{2}&-\eta_{2}&-H_{3}&(\pi\xi_{1},\cdots,\pi\xi_{i-1})^{\top}\\ -\mu_{1}&\boxed{1-P_{ii}}&-\mu_{2}&\boxed{-P_{ji}}&-\mu_{3}&\pi\xi_{i}\\ -H_{4}&-\eta_{3}&I_{j-i-1}-H_{5}&-\eta_{4}&-H_{6}&(\pi\xi_{i+1},\cdots,\pi\xi_{j-1})^{\top}\\ -\mu_{4}&\boxed{-P_{ij}}&-\mu_{5}&\boxed{1-P_{jj}}&-\mu_{6}&\pi\xi_{j}\\ -H_{7}&-\eta_{5}&-H_{8}&-\eta_{6}&I_{n-j}-H_{9}&(\pi\xi_{j+1},\cdots,\pi\xi_{n})^{\top}\end{array}\right].$
As the last column of Eq. (A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the $(i+1)$st row to make the $(i+1)$st row a zero vector. After dropping the $(i+1)$st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
$\left[\begin{array}[]{cc | c}1&\vec{1}_{n-1}^{\top}&0\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$ (A.18)
We multiply the matrix of Eq. (A.7) to the left side of Eq. (A.18) to get:
$\left[\begin{array}[]{cc | c}1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}&\vec{0}_{n-1}^{\top}&-\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}\\ -\alpha_{i}&I_{n-1}-Z_{i}&(\pi M_{i})^{\top}\end{array}\right].$
Therefore, by the first row of Eq. (A.19):
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}.$
Also, by the second row of Eq. (A.19):
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
and thus:
$\begin{array}[]{rcl}\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}&=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}+\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}\right]\\ &=&\left(I_{n-1}-Z_{i}\right)^{-1}\left[(\pi M_{i})^{\top}-\frac{\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}\left(\pi M_{i}\right)^{\top}}{1+\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\alpha_{i}}\alpha_{i}\right].\end{array}$
A12. Proof of Corollary 7
If $\pi M_{i}=\vec{0}_{n-1}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ according to Eqs. (17) and (18). Thus, $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$.
Conversely, if $\frac{\mathrm{d}\pi}{\mathrm{d}P_{ii}}=\vec{0}_{n}^{\top}$, then $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ and $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$. Applying $\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}=0$ to Eq. (17), we have:
$\vec{1}^{\top}_{n-1}(I_{n-1}-Z_{i})^{-1}(\pi M_{i})^{\top}=0.$ (A.20)
We plug Eq. (A.20) into Eq. (18) and use $\frac{\mathrm{d}\pi_{-i}}{\mathrm{d}P_{ii}}=\vec{0}_{n-1}$ to get:
$\left(I_{n-1}-Z_{i}\right)^{-1}\left(\pi M_{i}\right)^{\top}=\vec{0}_{n-1}.$
Therefore, $\left(\pi M_{i}\right)^{\top}=\left(I_{n-1}-Z_{i}\right)\vec{0}_{n-1}=\vec{0}_{n-1}$ and $\pi M_{i}=\vec{0}_{n-1}^{\top}.$
A13. Proof of Corollary 8
Since
[…]
$\pi M_{i}=\vec{0}_{n-1}^{\top}$ implies that $-(1-P_{ii})\pi M$, i.e.,
$\left(\pi_{1}\lambda_{i1}P_{i1},\cdots,\pi_{i-1}\lambda_{i,i-1}P_{i,i-1},-\pi_{i}(1-P_{ii}),\pi_{i+1}\lambda_{i,i+1}P_{i,i+1},\cdots,\pi_{n}\lambda_{in}P_{in}\right)\left[\begin{array}[]{c}\sigma_{1i}\\ \vdots\\ \sigma_{ni}\end{array}\right],$
is a zero vector except for its $i$th element. Thus, for any $t\neq i$, the zero value in the $t$th element implies that:
$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}P_{it}.$
We sum the above equations over all $t\neq i$ to get:
$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}\sum_{t\neq i}P_{it},$
which leads to:
$\sum\limits_{j\neq i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.21)
and:
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\odot P\right]_{ji}^{\top}=\pi_{i}.$
When $i$ runs all numbers in $\mathcal{N}$, thus, $\pi\left[\Lambda\odot P\right]^{\top}=\pi$.
Finally, if $\lambda_{ij}=\pi_{i}/\pi_{j}$, then Eq. (A.21) is clearly valid and so is Eq. (19). If $\lambda_{ij}=P_{ji}/P_{ij}$, then $\left[\Lambda\odot P\right]^{\top}=P$ and, therefore, Eq. (19) also holds.